### v1

# Magnetic Flux Emergence in Binary Neutron Star Remnants

Abstract
Using high-resolution AthenaK simulations of a twisted toroidal flux tube, we study the flux emergence of magnetic structures in the shear layer of a hot massive neutron star typical of a binary neutron star remnant. High-resolution simulations demonstrate that magnetic buoyant instabilities allow for emergence only for extremely large magnetic fields significantly exceeding $10^{17}~{}\mathrm{G}$, and more typical fields around $10^{16}~{}\mathrm{G}$ are instead dominated by hydrodynamic effects. Because merger remnants tend to be stable against hydrodynamic convection, our work places strong limitations on the mechanisms by which massive binary neutron star remnants can produce the magnetically-driven outflows needed to power jets.
I Introduction
During a binary neutron star (BNS) merger, the shearing motion between the stellar surfaces as they contact triggers turbulent instabilities, which help amplify the magnetic field to magnetar strengths within the first few milliseconds after merger [1, 2, 3, 4, 5, 6]. In the post-merger phase, this shear layer where the KHI is active settles to a region ${\sim}1~{}\mathrm{km}$ below the surface [7, 8], which is where the strongest fields are focused [5, 6]. Recent simulations suggest these strongly amplified fields can break out from the remnant star on relatively short timescales and help power short gamma-ray bursts (sGRBs) [9, 10]. However, it is not immediately clear how this breakout occurs; BNS remnants are typically stable against convection [7, 8, 11, 12], so this breakout cannot be purely hydrodynamical in nature.
A similar problem occurs in solar physics. Features such as coronal loops and prominences are direct consequences of strong magnetic fields emerging from the solar surface. These fields are generated by the solar dynamo, and they are propelled upward through the convective zone. Their progress then stalls inside the photosphere, which is stable against convection. A rather successful model for achieving emergence relies on magnetic buoyancy: as magnetic flux accumulates near the base of the photosphere, the magnetized region eventually becomes buoyant and accelerates upward [13].
To illustrate how this occurs, consider a pocket of fluid with a magnetic field $\mathbf{B}$ in pressure equilibrium with an ambient unmagnetized fluid. Therefore, the pressure in the pocket, $P^{\ast}$, is $\Delta P=B^{2}/8\pi$ lower than the ambient pressure, $P_{0}$. Assuming the pocket is either isentropic or in thermal equilibrium with the ambient fluid, the density in the pocket will be lower than the ambient fluid and thus experience a buoyant force.
A similar mechanism might be responsible for strong post-merger fields, which emerge from a BNS and create magnetically-driven outflows that could potentially power relativistic jets. A strong toroidal magnetic field develops inside the shear layer (see, e.g., Refs. [14, 15]), which could become buoyant as it settles into hydrostatic equilibrium with the surrounding fluid. From the Newtonian ideal magnetohydrodynamics (MHD) equations, the velocity $\mathbf{v}$ of the flux tube evolves as
$\rho\frac{D\mathbf{v}}{Dt}=-\nabla\left(P+\frac{B^{2}}{8\pi}\right)+\frac{1}{4\pi}\left(\mathbf{B}\cdot\nabla\right)\mathbf{B}+\rho\mathbf{g},$ (1)
where $D/Dt=\partial_{t}+\mathbf{v}\cdot\nabla$ is the advective derivative, $P$ is the pressure, $\rho$ is the density, $\mathbf{B}$ is the magnetic field, and $\mathbf{g}$ is the local gravitational acceleration. Consider an ambient unmagnetized fluid with state $\rho=\rho_{0}$ and $P=P_{0}$ in hydrostatic equilibrium. If we assume there exists a magnetic flux tube with $\mathbf{B}\neq 0$, pressure $P^{\ast}$ satisfying $P_{0}=P^{\ast}+B^{2}/8\pi$ (i.e., pressure equilibrium), and density $\rho^{\ast}$, the velocity is described by
$\rho^{\ast}\frac{D\mathbf{v}}{Dt}=\left(\mathbf{B}\cdot\nabla\right)\mathbf{B}-\Delta\rho\mathbf{g},$ (2)
where $\Delta\rho=\rho_{0}-\rho^{\ast}$. For a spherically symmetric neutron star, we can specialize to spherical coordinates and the $r$-direction. Therefore,
$\begin{split}\rho^{\ast}\frac{Dv^{r}}{Dt}=&\frac{1}{4\pi}\left(B^{r}\partial_{r}+B^{\theta}\partial_{\theta}+B^{\phi}\partial_{\phi}\right)B^{r}\\ &+\frac{\left(B^{r}\right)^{2}-B^{2}}{4\pi r}+\Delta\rho g(r),\end{split}$ (3)
where we have defined $\mathbf{g}=-g(r)\mathbf{\hat{r}}$ for convenience. If we consider $B^{\phi}$ to be the only non-zero component of the field (i.e., purely toroidal), then Eq. 3 simplifies to
$\frac{Dv^{r}}{Dt}=-\frac{B^{2}}{4\pi\rho^{\ast}r}+\frac{\Delta\rho}{\rho^{\ast}}g(r).$ (4)
Note the presence of the tension force $T\propto B^{2}/r$ on the right-hand side, which arises because we consider a toroidal field. This differs slightly from buoyancy in solar physics, which typically concerns relatively small sections of magnetized flux tubes. In the solar regime, a parallel-plane approximation is appropriate (e.g., [16, 17, 18, 13, 19, 20]), so there is no tension force at all. In the case of a BNS remnant, however, we are interested in the large-scale circumstellar toroidal field, and tension cannot be neglected. Recalling that $g(r)=Gm(r)/r^{2}$ for a spherically symmetric star, this implies that the field is only buoyant if
$B^{2}<4\pi\Delta\rho\frac{Gm(r)}{r}.$ (5)
This appears to impose a maximum field strength on $\mathbf{B}$, but the situation is more subtle than it appears because $\mathbf{B}$ is indirectly related to $\Delta\rho$ via $\Delta P$. It is therefore more appropriate to think of it as a condition on the equation of state (EOS) and permissible thermodynamic processes. Since $\Delta P=B^{2}/8\pi$, this implies the condition
$2\Delta P<\Delta\rho\frac{Gm(r)}{r}.$ (6)
Therefore, what matters is not the strength of the magnetic field, but how density changes during the thermodynamic processes that restore pressure equilibrium; the magnetic field is only indirectly constrained through this relationship, and it is usually a subdominant effect. For example, if we assume a general equation of state and an isentropic flux tube, this relation becomes (to first order)
$2\Gamma P_{0}\lesssim\rho_{0}\frac{Gm(r)}{r},$ (7)
where $\Gamma=(\partial\log P/\partial\log\rho)_{s}$ is the (not necessarily fixed) adiabatic index. Thus to first order, the strength of the magnetic field is not constrained, and whether or not the magnetized fluid is buoyant instead depends on the properties of the remnant. For a typical remnant neutron star with mass $M\sim 2~{}\textrm{M}_{\odot}$, radius $R\sim 10~{}\mathrm{km}$, and density $\rho\sim 10^{14}~{}\mathrm{g}/\mathrm{cm}^{3}$, this implies a maximum pressure of $P_{0}\sim 10^{34}~{}\mathrm{Ba}$, which is roughly compatible with realistic merger conditions with $P_{0}\sim 10^{33}~{}\mathrm{Ba}$, though less massive remnants may not be compact enough to satisfy this requirement. However, a semi-relativistic treatment (i.e., strong gravity and large internal energy but $v/c\ll 1$, see Appendix B) leads to a modified buoyancy condition which is usually less restrictive.
Note that this condition only determines buoyancy at a given moment and does not necessarily mean a toroidal field inside a BNS remnant will emerge; similar to classic hydrodynamic buoyancy, a stably stratified atmosphere can impede magnetic buoyancy, and finite-size effects (e.g., drag) as the scale height shrinks can also affect buoyancy [21, 16]. Nevertheless, this condition suggests that magnetic buoyancy may be a possible mechanism for embedded magnetic fields in BNS remnants to emerge.
Assuming the embedded magnetic field can emerge, the natural next step is to determine the correct time scale. Though we can estimate an evolution time scale from Eq. 4, it will be inaccurate for the length scale of interest because of large changes in the scale height and the stratification of the remnant. A better way to determine the time scale is to make some assumptions about the relationship between $B^{2}$, $\rho^{\ast}$, and $P^{\ast}$ and integrate Eq. 4 (or a relativistic equivalent) over the neutron star profile. However, these time scales are typically inaccurate; a time scale which neglects finite-size effects (e.g., drag forces) will be a significant underestimate, and approximating these effects requires additional assumptions which are not necessarily correct. Therefore, the best way to estimate the emergence time scale for a magnetically buoyant pocket in a BNS remnant is via full general relativistic magnetohydrodynamics (GRMHD).
In this paper, we consider an idealized model of a single flux tube embedded in the shear layer with a magnetic field strength typical of a post-merger BNS remnant using high-resolution (GRMHD) simulations. We begin by presenting our numerical setup in Sec. II. We then present our results in Sec. III, which demonstrate that the flux tube only emerges for extremely strong magnetic fields well above $10^{17}~{}\textrm{G}$, and the dynamics of tubes with more realistic fields around $10^{16}~{}\textrm{G}$ are dominated by hydrodynamic effects. Next, we discuss the implications of our work for magnetically driven outflows in realistic post-merger scenarios in Sec. IV, and we finally summarize our findings and suggest avenues for future research in Sec. V.
II Methods
II.1 Initial Data
We construct our model based on the results of Ref. [11]. As shown in Figures 3, 7, and 8 of Ref. [11], the remnant settles into a quasi-steady state approximately $10~{}\mathrm{ms}$ after merger. We perform nonlinear fits as functions of density (see Appendix A) to the temperature and lepton fraction profiles in order to compute a 1D finite-temperature equilibrium slice of the HS(DD2) EOS (hereafter DD2)[22, 23], which we then use to generate a static neutron star via the Tolman-Oppenheimer-Volkoff equations [24, 25]. Using a central density $\rho=2.2\times 10^{-3}~{}\textrm{M}_{\odot}^{-2}\approx 1.36\times 10^{15}~{}\mathrm{g}/\mathrm{cm}^{3}$, the resulting star has a mass of $M\approx 2.43~{}\textrm{M}_{\odot}$ and a radius $R\approx 8.35~{}\textrm{M}_{\odot}=12.3~{}\mathrm{km}$. Because our table only covers a finite density range, we define the edge of the star as the point dropping below $\rho=10^{-5}~{}\textrm{M}_{\odot}^{-2}\approx 6.18\times 10^{13}~{}\mathrm{g}/\mathrm{cm}^{3}$.
Because we are interested in the dynamics of buoyant magnetic structures, we add a toroidal flux tube of radius $s_{0}$ satisfying
$\sqrt{\gamma}B^{\phi}$ $=B_{0}e^{-s^{2}/\sigma^{2}}$ (8a)
$\sqrt{\gamma}B^{\psi}$ $=qsB_{0}e^{-s^{2}/\sigma^{2}},$ (8b)
where $s^{2}=r^{2}+R^{2}-2Rr\sin{\psi}$ is the coordinate distance from a ring of radius $R$ in the coordinate equatorial plane, $\psi$ is the poloidal angle of the torus, $q$ is a twist parameter, and $\gamma$ is the determinant of the spatial metric $\gamma_{ij}$. While this magnetic field trivially satisfies the divergence-free condition in toroidal coordinates, $B^{r}$ and $B^{\theta}$ will both depend on $r$ and $\theta$. Therefore, to limit numerical errors in the divergence-free condition during the coordinate transformation, we instead compute $B^{r}$ and $B^{\theta}$ from the azimuthal component of the vector potential, which is the same in both spherical and toroidal coordinates (note that other contributions vanish due to axisymmetry):
$A_{\phi}=\begin{cases}-\frac{1}{2}qB_{0}\sigma^{2}e^{-s^{2}/\sigma^{2}}&s\leq s_{0}\\ -\frac{1}{2}qB_{0}\sigma^{2}e^{-s_{0}^{2}/\sigma^{2}}&s>s_{0}.\end{cases}.$ (9)
A realistic magnetic field in a merger is not completely axisymmetric, which can trigger additional magnetic buoyant instabilities beyond the simple form of magnetic buoyancy already discussed. Among the most important of these effects is the so-called Parker instability (or undular instability) [26, 27]. Consider a longitudinal perturbation (i.e., parallel to the field lines) to the density or velocity to a state otherwise in equilibrium. The field lines will bend and form a series of peaks and troughs. The fluid will attempt to slide along the field lines, which causes the peaks to become lighter and more buoyant while the troughs become heavier and thus fall. The Parker instability can often grow in a stratified medium where other forms of magnetic buoyancy cannot, which makes it an attractive prospect for enhancing buoyancy in our strongly stratified remnant profile.
Consequently, we also add a small density perturbation of the form
$\rho=\rho_{0}\left(1-A\cos\left(k\phi\right)\right)$ (10)
along the length of the tube where $A$ is a specified amplitude and $k$ is an integer determining the harmonic; because we keep the pressure fixed, this indirectly adds a perturbation to the temperature inside the flux tube. These longitudinal perturbations add a small amount of hydrodynamic buoyancy to the system, causing the field lines to bend and possibly triggering the Parker instability.
After constructing the flux tube, we enforce mechanical equilibrium by setting the tube’s pressure to $P^{\ast}=P_{0}-P_{B}$, where $P_{0}$ is the ambient fluid pressure and $P_{B}$ is magnetic pressure ($P_{B}=b^{2}/8\pi$ in CGS units, where $b^{\mu}$ is the magnetic field in the fluid’s rest frame). We then reset the fluid density to enforce thermal equilibrium with the (perturbed) ambient medium to create a buoyant force. In a true post-merger scenario, the thermal diffusion time scale is very long compared to the dynamical time scale, so calculating the density difference assuming an adiabatic process rather than an isothermal one may produce a more realistic flux tube. However, one can show that enforcing thermal equilibrium results in a larger density difference than an isentropic system, independent of the EOS, because $(\partial\rho/\partial P)_{T}>(\partial\rho/\partial P)_{s}$ is equivalent to requiring the heat capacities to satisfy $C_{P}/C_{V}>1$. Therefore, these tests tend to overestimate how buoyant a flux tube is and provide something of a best-case scenario for emergence.
Table 1: A summary of the simulations performed, including the initial maximum field strength, the radius $s_{0}$ of the flux tube, the estimated emergence time, and the estimated velocity at emergence.
Name $\sqrt{b^{2}_{\mathrm{max}}}$ (G) $A$ $s_{0}$ ($\textrm{M}_{\odot}$) $t_{\mathrm{em}}$ (ms) $v_{\mathrm{em}}$
H7e17_5 $7.540\times 10^{17}$ $0.05$ $0.25$ 0.216 0.038
H7e17_5_B $7.540\times 10^{17}$ $0.05$ $1.0$ 0.201 0.013
H7e17_0 $7.540\times 10^{17}$ 00 $0.25$ 0.224 0.034
H1e17_5 $1.257\times 10^{17}$ $0.05$ $0.25$ N/A N/A
H1e17_0 $1.257\times 10^{17}$ 00 $0.25$ N/A N/A
H1e16_5 $1.257\times 10^{16}$ $0.05$ $0.25$ N/A N/A
H6e15_5 $6.283\times 10^{15}$ $0.05$ $0.25$ N/A N/A
H0_5 00 $0.05$ $0.25$ N/A N/A
For most of our runs, we consider a flux tube of inner radius $s_{0}=0.25~{}\textrm{M}_{\odot}\approx 369~{}\mathrm{m}$ and an outer radius $R$ matching the position of the peak temperature ($r=7.114~{}\textrm{M}_{\odot}$), which we take as a proxy for the position of the shear layer. We additionally set the width parameter $\sigma=0.3$ and add a small twist with $q=1$ to increase rigidity of the flux tube. The perturbation wavenumber is set to $k=18$.
Across our runs we vary the magnetic field strength and the perturbation amplitude. Based on the relativistic generalization of Eq. 5 (Eq. 24), only very weak fields ($B\lesssim 4\times 10^{12}~{}\textrm{G}$) are completely incapable of buoyancy. Due to the finite width of the flux tube and discretization effects, the necessary field strength for buoyancy can deviate somewhat from this ideal condition, but it serves as a useful guideline nonetheless. We test fields at $7.540\times 10^{17}~{}\textrm{G}$, $1.257\times 10^{17}~{}\textrm{G}$, $1.257\times 10^{16}~{}\textrm{G}$, and $6.283\times 10^{15}~{}\textrm{G}$. For every field strength, we perform a run with a perturbation amplitude $A=0.05$. We perform an additional test with $A=0$ for some cases to help investigate the role that the Parker instability plays in the perturbed runs. As a control, we also perform a run with no magnetic field but with an $A=0.05$ density perturbation; this will help disentangle the effects of ordinary hydrodynamic buoyancy from those caused by the Parker instability. To help track the evolution of the tube for this case, a passive tracer scalar $s$ is added with $s=1$ inside the tube and $s=0$ everywhere else.
Each run is recorded in Table 1, with the names following the format “Px_A”. “P” is a letter indicating the stellar profile, “x” indicates the magnetic field with the leading digit and order of magnitude, and “A” is the strength of the amplitude in percent. Thus the run with a $7.540\times 10^{17}~{}\textrm{G}$ field embedded in the hot remnant profile with an $A=0.05$ perturbation is labeled “H7e17_5“, while “H0_5” corresponds to the base case with a density perturbation but no magnetic field.
Since the shear layer is ${\sim}3~{}\mathrm{km}$ wide in a realistic BNS merger [8], a flux tube of $s_{0}=0.25~{}\textrm{M}_{\odot}$ may be much smaller in size than the toroidal field structure which develops in a realistic merger simulation. Because we may approximate the field as a superposition of weakly twisted flux tubes, tests with small flux tubes should be qualitatively correct when determining if a particular flux tube can emerge. Nevertheless, we do include one additional setup with the H7e17_5 configuration with a larger flux tube as well. For this run, we set $s_{0}=1.0~{}\textrm{M}_{\odot}\approx 1.48~{}\mathrm{km}$ and reduce the twist coefficient to $q=0.25$; since the twist scales linearly with the tube radius $s$, this is necessary to ensure the magnetic field remains primarily toroidal. We label this run H7e17_5_B to distinguish it from the case with a small flux tube.
We note that our strongest fields are unrealistically large for a BNS post-merger; high-resolution BNS simulations starting with realistic initial field strengths do not typically produce fields with peak strengths much in excess of $10^{17}~{}\textrm{G}$, and average strengths are typically ${\sim}10^{16}~{}\textrm{G}$ [3, 4, 28, 5, 14, 15]. Though no full-scale merger simulation has sufficient resolution to capture all scales of magnetic turbulence, local simulations confirm these findings [29, 30].
II.2 Numerical Methods
We evolve the star with the GRMHD code AthenaK [31, 32, 33], which uses a second-order finite-volume method to evolve the fluid and upwind constrained transport to ensure that $\nabla\cdot\mathbf{B}=0$. For these tests, we choose WENOZ as our reconstruction scheme [34] and enable a first-order flux correction (FOFC) method [35]. The FOFC scheme greatly improves the stability of the evolution near the surface of the neutron star, particularly when the magnetization is high. Because we assume that the remnant has achieved steady state, we fix the spacetime to the initial TOV solution.
Our computational domain is a wedge in the equatorial plane covering $r\in[6~{}\textrm{M}_{\odot},24~{}\textrm{M}_{\odot}]$, $\theta\in[\pi/2-\pi/9,\pi/2+\pi/9]$, and $\phi\in[-\pi/9,\pi/9]$, which amounts to full two oscillations for the $k=18$ perturbation described in Sec. II.1. We use a base domain of $600\times 180\times 180$ cells with a single refinement region over $r\in[6~{}\textrm{M}_{\odot},12\textrm{M}_{\odot}]$ for an effective radial resolution of $\Delta r=0.015~{}\textrm{M}_{\odot}\approx 22~{}\mathrm{m}$ inside the star. The angular resolution is sufficient that $r\Delta\theta\approx r\sin{\theta}\Delta\phi\approx\Delta r$ over the region of interest from approximately $7~{}\textrm{M}_{\odot}$ to $9~{}\textrm{M}_{\odot}$.
The upper boundary of $r$ is a “diode” outflow boundary (the radial momentum must be nonnegative), and the $\theta$ and $\phi$ boundaries are periodic. For the inner boundary of $r$, we keep the rest-mass density $\rho$ and pressure $P$ fixed, but we apply a reflecting boundary to the magnetic field $B^{i}$ and the radial fluid momentum density $\alpha\sqrt{\gamma}\rho Wv^{r}$, where $\gamma$ is the determinant of the spatial metric, $\alpha$ is the lapse, $v^{r}$ is the radial component of the fluid three-velocity, and $W=(1-v^{2})^{-1/2}$ is the Lorentz factor. This condition ensures that the density flux across the boundary is approximately zero. The $Wv^{\phi}$ and $Wv^{\theta}$ velocity components are also reflected across the boundary with the parity inverted. Though these components are parallel to the boundary, we found that inverting the parity improves stability; otherwise, small numerical errors can accumulate in the velocity, and a simple reflection can amplify these errors in a way that gives the star an unphysical rotation, while flipping the sign of the velocity components damps them.
This work represents the first use of 3D EOS tables in AthenaK, which are implemented following the same prescription as in GR-Athena++ [36]. To approximate their effects on the remnant, we modify the full DD2 EOS table with the contributions of trapped neutrinos following the prescription in Ref. [37]. Therefore, the evolved species fraction is now the (electron) lepton fraction, $Y_{l}=Y_{e}+(n_{\nu_{e}}-n_{\nu_{\overline{e}}})/n_{b}$, which is advected along with the fluid flow. As $Y_{e}$ becomes a function of $Y_{l}$, $n_{b}$, and the temperature $T$, this prescription allows us to approximate the effects of trapped, fast-equilibriating neutrinos without needing to evolve them directly.
We evolve each simulation either until shortly after the flux tube emerges or until it becomes clear that the flux tube will not emerge; the exact length depends on the initial conditions and the observed dynamics.
III Results
We begin by discussing the non-magnetized base case, H0_5. As shown in Fig. 1, the density perturbation causes lighter parts of the tracer region to rise while the overdense regions sink, in agreement with standard hydrodynamic buoyancy. However, because the remnant is stable against convection, the perturbations slow down and eventually settle into layers of the remnant with similar densities by $t=0.39~{}\textrm{ms}$.
Figure 1: A slice of the $r$-$\phi$ plane at $t=0.39~{}\textrm{ms}$ comparing $|B|$ from H1e16_5 (left) to the tracer scalar $s_{1}$ from H0_5 (right). The dashed black line marks the surface of the star, and the dot-dashed green lines mark the boundaries of the initial position of the flux tube. The nearly identical evolution indicates that the dynamics of H1e16_5 are driven primarily by hydrodynamic buoyancy from the original perturbation rather than magnetic buoyancy.
III.1 Effects of magnetic field strength
The H6e15_5 and H1e16_5 cases, which feature the weakest magnetic fields, are qualitatively similar to the H0_5 case. From comparing the magnetic field of H1e16_5 to the tracer of H0_5 in Fig. 1, it is clear that the fluid within the flux tube rises and falls exactly the same way as predicted by hydrodynamic buoyancy alone. This is not surprising, as the magnetization $\sigma_{b}=b^{2}/4\pi\rho c^{2}$ is initially ${\sim}10^{-5}\ll\Delta\rho/\rho$. The magnetic field does appear to become more concentrated in between the peaks and troughs, increasing by a factor of ${\sim}5$ in this brief time.
To determine if this is dynamically important, we examine the energetics of the system. We start with the stress-energy tensor for ideal MHD,
$T^{\mu\nu}=(\rho h+b^{2})u^{\mu}u^{\nu}+\left(P+\frac{b^{2}}{2}\right)g^{\mu\nu}-b^{\mu}b^{\nu},$ (11)
where $\rho$ is the rest-mass density, $h$ is the total specific enthalpy, $b^{\mu}$ is the magnetic field in the fluid frame, $P$ is the fluid pressure, and $u^{\mu}$ is the four-velocity. Also note that we have switched from the Gaussian convention for electromagnetism to Lorentz-Heaviside to better align with the form of the GRMHD equations as used in evolution codes. In the Valencia formalism used in AthenaK, we evolve the energy as $\tau=\alpha^{2}T^{00}-\rho W$, where $\alpha$ is the lapse and $W$ is the Lorentz factor, which is equivalent to
$\tau=\rho W\left(hW-1\right)+B^{2}-P-\frac{b^{2}}{2}.$ (12)
It is easy to verify in the Newtonian limit that this reduces to the usual Newtonian expression for the total non-gravitational energy in MHD. This expression has contributions from the internal, kinetic, and magnetic energy of the fluid. Similar to Ref. [38], we express the kinetic energy $E_{K}$ and magnetic energy $E_{B}$ as
$E_{K}$ $=\int\rho W\left(W-1\right)\mathrm{d}V,$ (13)
$E_{B}$ $=\int\left(B^{2}-\frac{b^{2}}{2}\right)\mathrm{d}V,$ (14)
where $dV$ includes the appropriate volume element. Since $\tau$ does not include the gravitational energy of the system, it is not conserved. However, $-\alpha(T^{0}_{0}+\rho u^{0})$ does include these contributions and is therefore conserved in stationary spacetimes. It is easy to verify that $\alpha T^{0}_{0}+\alpha\rho u^{0}+\tau\approx-\rho\Phi$ in the Newtonian limit, where $\Phi$ is the gravitational potential. We therefore define
$E_{\mathrm{tot}}=-\int\left(T^{0}_{0}+\rho u^{0}\right)\alpha\mathrm{d}V,$ (15)
which we use to normalize the energy of our system. We caution that $E_{\mathrm{tot}}$ is not precisely invariant in due to boundary effects and truncation error from evolving $\tau$ instead of $\alpha(T^{0}_{0}+\rho u^{0})$. However, we have confirmed that our results are the same if we normalize by $E(t=0)$ or $E(t)$, indicating that conservation violations are insignificant for our analysis.
Figure 2: The magnetic energy $E_{B}$ (solid lines) and kinetic energy $E_{K}$ (dashed lines) as a function of time normalized by the total energy for all the runs with perturbations. The dotted vertical line marks the emergence time for H7e17_5.
In Fig 2, we see that $E_{B}\ll E_{K}$ for both H6e15_5 and H1e16_5, and $E_{K}$ deviates only slightly from H0_5 after ${\sim}0.3~{}\textrm{ms}$. The magnetic energy grows steadily until ${\sim}0.5~{}\textrm{ms}$, then begins to decay with $E_{K}$. The simultaneous decay of both these quantities suggests that both systems are stable and are converging toward a new equilibrium.
The H1e17_5 almost immediately shows deviations from the hydrodynamic case. Due to the increased magnetic tension relative to the perturbation, the flux tube does not bend as readily, and its peaks rise only a short distance before stalling and oscillating in place. Fig. 3 shows the flux tube at both $t=0.39~{}\textrm{ms}$ and $t=1.31~{}\textrm{ms}$ to demonstrate this; though there is some expansion and diffusion apparent in the flux tube at later times, the location of the core is nearly unchanged.
Figure 3: Snapshots of $\tilde{B}^{\phi}$ in the $r$-$\phi$ plane from H1e17_5 at $t=0.39~{}\textrm{ms}$ (left) and $t=1.31~{}\textrm{ms}$ (right). The location of the flux tube is roughly the same in both, but some expansion and diffusion is apparent at the later time. The dashed black line indicates the location of the stellar surface from the TOV solution, and the dot-dashed green lines mark the boundaries of the flux tube at $t=0$.
Fig. 2 corroborates this picture; for H1e17_5, $E_{K}$ peaks at a lower energy and begins to decay sooner, indicating that this case is also stable. The magnetic energy decays as well, albeit more slowly than the kinetic energy. If we consider a volume $V=A\ell$, where $A$ is some changing cross-sectional area and $\ell$ is an approximately fixed length, flux conservation implies that $E_{B}\approx B^{2}V\propto A^{-1}$, so we attribute this slow decay in magnetic energy to the slow diffusion or expansion of the flux tube as demonstrated in Fig. 3.
Figure 4: The magnitude of the magnetic field $|B|$ in the $r$-$\theta$ plane from H7e17_1_5 at $t=0~{}\textrm{ms}$ (top-left), $t=0.15~{}\textrm{ms}$ (top-right), $t=0.25~{}\textrm{ms}$ (bottom-left) and $t=0.34~{}\textrm{ms}$ (bottom-right). The contours in black, blue, and red denote $10^{15}~{}\textrm{G}$, $10^{16}~{}\textrm{G}$, and $10^{17}~{}\textrm{G}$, respectively. As the flux tube rises, it expands into the umbrella shape characteristic of a stratified environment with a decreasing scale height. The dashed black line marks the location of the stellar surface from the TOV solution.
For H7e17_5, which has the strongest magnetic field, the flux tube successfully emerges. This emergence process is shown in Figs. 4 and 5. The flux tube stretches out into a characteristic umbrella shape due to the decreasing scale height as it rises, similar to what occurs in solar studies (e.g., Fig. 4 of Ref. [39]). As it reaches the surface, the flux tube rapidly expands and pushes out a wave of material in front of it (see Fig. 6).
Figure 5: Streamlines of the magnetic field for H7e17_5 (top) and H1e17_5 (bottom) at the initial time (left) and a late time (right). The surface contour marks $\rho=10^{-5}~{}\textrm{M}_{\odot}^{-2}\approx 6.2\times 10^{12}\mathrm{g}/\mathrm{cm}^{3}$, which is the approximate location of the surface. Note that the color scale is normalized to the maximum field strength in each run. H7e17_5 expands rapidly as it emerges, while H1e17_5 remains trapped beneath the surface and largely intact.
Figure 6: A comparison of $|B|$ (top) and $\rho$ (bottom) in the $r$-$\phi$ plane from H7e17_5 at $t=0.34~{}\textrm{ms}$. Much of the flux tube remains near the surface, but the wave of material pushed out by it is also visible. The dashed black line marks the location of the stellar surface from the TOV solution, and the dot-dashed green lines mark the boundaries of the flux tube at $t=0$.
To measure the emergence time and properties of the entrained material, we integrate the mass flux $\dot{M}$ across surfaces at $r=8.35~{}\textrm{M}_{\odot}$, $r=8.4~{}\textrm{M}_{\odot}$, $r=8.5~{}\textrm{M}_{\odot}$, $r=8.6~{}\textrm{M}_{\odot}$, and $r=8.7~{}\textrm{M}_{\odot}$. We define the emergence time as the peak mass flux $\dot{M}$ across the surface at $r=8.35~{}\textrm{M}_{\odot}$, which is the approximate location of the remnant’s surface. By looking for where $\dot{M}$ peaks in the other integration surfaces, we can measure the approximate speed of the flux tube at emergence, which we do by fitting a quadratic polynomial for $r(t)$ to these peak times and taking its time derivative.
Fig. 7 shows the time-integrated mass across each integration surface along with the estimated emergence time. The flux tube emerges at $t_{\mathrm{em}}\approx 0.216~{}\textrm{ms}$ for H7e17_5 with a speed of $v_{\mathrm{em}}\approx 0.038c$. The total entrained mass peaks around $M\approx 1.5\times 10^{-3}~{}\textrm{M}_{\odot}$ shortly after emergence, but it rapidly declines as material falls back onto the remnant. The mass across each integration surface also rapidly decreases with distance, suggesting that little to no mass is unbound [^1] A better way to estimate unbound ejecta is using the geodesic and/or Bernoulli criteria, but this must be done farther away from the remnant to be accurate. Because our simulations only span $[-\pi/9,\pi/9]$ in the $\theta$ direction with periodic boundary conditions, the emerging fluid begins to interact with itself not long after passing the surface at $r=8.7~{}\textrm{M}_{\odot}$. Therefore, the dynamics of the fluid will not be reliable at a radius where a proper ejecta calculation would be meaningful..
Figure 7: Total integrated mass flux for H7e17_5 at high (solid lines) and medium resolution (dotted lines). The integrated flux is also shown for H7e17_0 at high resolution (dash-dotted line). The low-resolution data for H7e17_5 is not shown because it does not emerge. Black vertical lines indicate emergence time as estimated by the peak instantaneous flux across the first integration surface at $r=8.35~{}\textrm{M}_{\odot}$.
Unsurprisingly, Fig. 2 demonstrates that E7e17_5 has larger $E_{B}$ and $E_{K}$ than in the weak-field runs. $E_{K}$ follows the same qualitative trend as the other runs, with a rapid rise followed by a peak and slow decay in time (possibly with small oscillations). Unlike the other runs, however, $E_{B}$ immediately decays, while it remains nearly constant in E1e17_5 and grows for some time in both E1e16_5 and E6e15_5. There is also a crossover between $E_{B}$ and $E_{K}$ at the same approximate time emergence occurs, though, as will be discussed later, this may simply be coincidence.
III.2 Effects of density perturbation
We now consider the effects of the perturbation by analyzing the non-perturbed cases, beginning with E1e17_0. Since E1e17_5, which has a density perturbation, does not emerge, it is not surprising that E1e17_0 does not, either, and instead rises only for a short time before stalling. As shown in Fig. 8, $E_{K}$ peaks at ${\sim}1/3$ its value in E1e17_5, and it rapidly decays by nearly an order of magnitude in ${\sim}0.3~{}\textrm{ms}$. Additionally, though the evolution of $E_{B}$ in E1e17_0 is initially similar to E1e17_5, it does not experience the same slow rise, but instead slowly decays over time.
Figure 8: Like Fig. 2, but comparing the cases with no perturbation to their perturbed counterparts. The dotted vertical lines mark the emergence times for H7e17_0 and H7e17_5.
Comparing the E7e17_0 case to E7e17_5, we find that the perturbation has very little effect on the dynamics of the system. This is consistent with our initial conditions, as the change in density from our equilibrium conditions on the flux tube ($\Delta\rho/\rho_{0}\approx 0.6$) is much larger than the change in density from the perturbation ($A=0.05$). According to Fig. 8, the presence of a perturbation does not affect the energetics of the system in any significant way. However, there is a small effect on both the emergence time and the mass pushed out by the emerging flux tube. Fig. 7 shows that emergence in E7e17_0 is slightly delayed to $t_{\mathrm{em}}\approx 0.224$ and slightly slower at $v_{\mathrm{em}}\approx 0.034c$ compared to E7e17_5, though the entrained mass is slightly larger. Nevertheless, the mass again rapidly decays with distance, suggesting that the amount of unbound ejecta is small.
III.3 Effects of flux tube size
Thanks to its extended profile, the H7e17_5_B run emerges at ${\sim}0.201~{}\textrm{ms}$ according to the mass flux criterion, slightly earlier than H7e17_5. However, we note that this time estimate requires ignoring the first peaks in the data and comparing with slice plots of the data. The core of the flux tube remains trapped, likely due to increased drag from the larger cross section, so the mass flux is smaller than H7e17_5. Because of the larger size of the flux tube, the integrated mass flux peaks later than H7e17_5 and is only ${\sim}20\%$ smaller. However, it is harder to distinguish the results from surface oscillations, as seen in the top half of Fig. 9. This contamination also affects the additional integration surfaces (not shown), so the velocity at emergence is significantly more uncertain. Nevertheless, we estimate it to be ${\sim}0.013c$, approximately a third of H7e17_5 and H7e17_0. The energetics for H7e17_5_B are shown in the bottom plot of Fig. 9. Qualitatively, the energetics are similar to H7e17_5, including the presence of a crossover around the time of emergence.
Figure 9: (Top) The integrated mass flux across the surface of the star for the resolution study with H7e17_5_B. (Bottom) Like Fig. 2, but comparing the different resolutions of H7e17_5_B. Dotted vertical lines in both plots mark the measured emergence time using the peak of $\dot{M}$.
III.4 Effects of resolution
To assess the effects of resolution, we include two additional tests for H7e17_5 at effective resolutions of ${\sim}79~{}\mathrm{m}$ (which we label “LR”) and ${\sim}44~{}\mathrm{m}$ (or “MR”) at the location of the flux tube, which correspond to roughly $9$ and $16$ cells across the flux tube, respectively. The standard run at ${\sim}22~{}\mathrm{m}$ is denoted as “HR” for these comparisons. The LR run does not emerge at all, as the flux tube expands horizontally too rapidly and stalls near the surface of the star. As shown in Fig. 7, the MR run features a slightly delayed emergence time with a smaller mass flux compared to the HR run reported in Table 1. Similar to the LR case, this appears to be due to the tube spreading more horizontally, which reduces how much of the flux tube fully emerges.
According to Fig. 10, the energetics are qualitatively similar between the different resolutions, though the lower resolutions see $E_{B}$ decay faster and less growth in $E_{K}$, consistent with additional diffusion in the flux tube. The crossover point between the two exists in all three runs, including the LR run that does not emerge, and seems to move earlier in time as the resolution decreases. As mentioned earlier, however, the MR run emerges slightly later than HR, and LR does not emerge at all. This suggests that the crossover time is not directly related to the emergence time, and the agreement between the two in the HR run is coincidence.
Figure 10: Like Fig. 2, but for the different resolutions of H7e17_5. Dotted vertical lines mark the emergence time.
Buoyancy should depend on resolving not only the flux tube but also the density and pressure scale heights, $H_{\rho}\equiv-(d\rho/dr)^{-1}\rho$ and $H_{P}\equiv-(dP/dr)^{-1}P$, respectively. Even if $H_{\rho}$ or $H_{P}$ are initially resolved, both scale heights will decrease as we approach the surface and become unresolved earlier and earlier as we decrease the resolution. We therefore also perform a resolution study for H7e17_5_B to assess the effects of under-resolving $H_{\rho}$ and $H_{P}$ more and more while the flux tube remains adequately resolved. In addition to LR and MR configurations, we also add a very low resolution (VLR) configuration with $\Delta r\approx 158~{}\mathrm{m}$ at the location of the flux tube. Therefore, the VLR, LR, MR, and HR runs have ${\sim}18$, ${\sim}37$, ${\sim}67$, and ${\sim}134$ cells across the flux tube. In all cases, part of the flux tube emerges successfully, though the core with the highest magnetization remains trapped. The mass flux increases monotonically with resolution except for the VLR case (see Fig. 9), which appears to push out noticeably more mass than any other resolution. However, as in H7e17_5, the integrated mass decreases rapidly across the other integration surfaces (not shown), even for VLR, suggesting that it is still bound.
There is little consistency in the merger time, and the effects of surface oscillations make it difficult to determine exactly where emergence occurs, particularly for the LR case. Though the flux tube is well resolved for the LR, MR, and HR cases (and arguably adequately resolved for the VLR case), it is likely that we are not in a completely convergent regime, even if we ignore the measured emergence times.
The energetics in Fig. 9 corroborate this picture; the behavior for $E_{B}$ and $E_{K}$ are not entirely consistent with resolution, particularly for the VLR case, which has a slightly larger initial $E_{B}$, and $E_{K}$ grows larger than both MR and LR before decaying more rapidly (which is expected). However, the qualitative dynamics are similar between all cases, and the predicted differences in emergence time, though noticeable on the scale of our simulations ($t\sim 1~{}\mathrm{ms}$ or less), would be small over the course of a full BNS merger.
IV Discussion
For our magnetic field configuration, we observe that only extremely large fields are capable of emerging. Because our field is a single twisted loop surrounding the entirety of the star, however, there is a strong magnetic tension force which, using Eq. 24, initially exceeded $50\%$ of the gravitational force on the flux tube in some of our tests. Alternate field configurations may be less susceptible to this tension force, but numerical simulations suggest that post-merger fields, independent of their initial (realistic) conditions, are dominated by a large-scale toroidal component [14, 40], which limits the possibilities for alternate field structures.
Even if this were not the case, this likely is only important for fields of ${\sim}10^{17}~{}\textrm{G}$; we observe that the evolution of the flux tube at weaker field strengths around $10^{16}~{}\textrm{G}$ are dominated by ordinary hydrodynamic buoyancy. Therefore, it seems unlikely that altering the field configuration is enough to help these features to emerge. We do not consider the effects of rotation in our model, but we remind the reader that rotation tends to inhibit convective processes, not enhance them [41, 42, 12].
This result may initially seem at odds with works suggesting that strong magnetically driven outflows may develop from a BNS remnant, such as Most and Quataert [9] and Combi and Siegel [10]. This is not necessarily the case; the demonstrated flare-like features and outflows in Most and Quataert [9] originate from strong magnetic fields exceeding $10^{17}~{}\textrm{G}$. Similarly, the post-merger jet in Combi and Siegel [10] develops in a remnant with an average toroidal field strength nearing and isolated pockets exceeding $10^{17}~{}\textrm{G}$. However, we again stress that such field strengths are atypically large compared to observation of magnetars and what is expected in realistic merger remnants. Even under these circumstances, such fields are not a guarantee that magnetically-driven outflows develop from the remnant. Musolino et al. [43], for example, suggest that, even with an extremely large magnetic field, these outflows may instead originate from the disk.
Similar to alternate configurations, buoyant instabilities like the Parker instability are likely only important for a small range of field strengths in excess of $10^{17}~{}\textrm{G}$. Despite a strong longitudinal perturbation, the Parker instability does not appear to develop in H1e17_5; rather, the flux tube oscillates. This agrees with results from Ref. [43], for instance, which suggest that the post-merger disk can be unstable to the Parker instability, but the remnant itself should be stable, even for a very large magnetic field. Larger fields like H7e17_5 may be Parker-unstable, but the emergence time scale is so short that the instability cannot develop quickly enough to be dynamically important (as seen when compared with H7e17_0). There likely exists some middle ground between these two states where the Parker instability develops quickly enough to accelerate the emergence time scale, but the required field strengths may not be achievable in a realistic merger setting.
As we consider cases which do emerge, it is interesting to note that whether or not emergence occurs seems to depend only mildly on resolution so long as the buoyant feature is adequately resolved. Though this is a necessary condition for emergence, one would not necessarily expect it to be sufficient; under-resolving $H_{\rho}$ could lead to an incorrect density difference between the ambient medium and the interior of the buoyant feature, therefore leading to an incorrect emergence time scale. Similarly, under-resolving $H_{P}$ could cause the buoyant feature to expand too much or too little as it rises, thus indirectly affecting the density difference and limiting buoyancy. In our runs, $H_{\rho}\approx 2.43~{}\mathrm{km}$ and $H_{P}\approx 827~{}\mathrm{km}$ at the shear layer, but both these quantities will rapidly decrease as the flux tube approaches the surface. Therefore, $H_{\rho}$ will only be moderately resolved at the beginning for the VLR run, and $H_{P}$ will only be resolved by the HR and possibly MR runs to start. Though we see some differences in the emergence time (which may not be entirely accurate for H7e17_5_B), they are small enough that they would be insignificant in the context of a full merger simulation. This suggests that the timescales shown in global BNS runs with $\Delta x\lesssim 180~{}\mathrm{m}$ are likely reasonable despite the limited resolution.
Secondary characteristics such as the amount of mass pushed out during emergence are more strongly affected by resolution. Though our evidence is suggestive that these outflows remain bound, it is not definitive. Furthermore, such outflows in a realistic merger would also carry some angular momentum and interact with the magnetic fields already existing outside the remnant. A scenario where some of this material becomes unbound is not inconceivable, though how strong of an effect this is would be determined best by high-resolution global simulations.
V Conclusion
In this paper, we have investigated the dynamics of magnetically buoyant flux tubes inside a neutron star remnant. We began by reviewing the concepts of magnetic buoyancy in a Newtonian context, though we have also derived a semi-relativistic buoyancy condition for a TOV spacetime as well. For a neutron star, a strong magnetic tension force appears due to the toroidal structure of the magnetic field. This tension force inhibits buoyancy and leads to an indirect condition on the equation of state for emergence to be possible, even in the absence of stable stratification. Relativistic enhancements to gravity often relax this constraint sufficiently for our remnant profile that flux tubes with realistic post-merger fields are all at least marginally buoyant. Nevertheless, we note that this tension force means that intuition from similar emergence problems in solar physics is limited.
While our model only uses a simple, idealized field configuration, it captures the essential features of a more realistic post-merger field. We additionally impose thermal equilibrium rather than adiabaticity on our initial data, which increases the buoyancy of the flux tube. Therefore, our results overestimate how buoyant a realistic post-merger field would be.
Nevertheless, we also observe that these flux tubes can only overcome the stably stratified profile of the remnant if the magnetic field is significantly more than $10^{17}~{}\textrm{G}$, which is often considered unphysically large. In agreement with Musolino et al. [43], this suggests that the strong magnetically-driven outflows needed for BNS remnants to power relativistic jets may originate in the disk rather than the remnant.
Though our work suggests emergence itself is somewhat independent of resolution so long as the large-scale magnetic field is well resolved, secondary features like the mass flux may be more sensitive to resolution. Therefore, a natural extension of this work is to explore emergence with high-resolution global simulations where potential outflows can interact with the remnant disk and flows. However, it is not immediately clear if such work would be useful; if a BNS merger with a realistic initial field strength cannot self-consistently develop a field well above $10^{17}~{}\textrm{G}$ or demonstrate emergence at lower field strengths, high-resolution simulations may be of little practical worth.
A possible study aimed toward probing emergence at lower field strengths could consider the effects of the EOS; DD2 is a so-called “stiff” EOS, where a small change in density produces a large shift in pressure, thus supporting heavier remnants with larger radii. Due to the higher compactness, intuitively one would expect flux emergence in a long-lived remnant with a softer EOS to be less likely because of increased magnetic tension. However, enforcing pressure equilibrium would also lead to a larger relative difference in density and enhance buoyancy, so it is not immediately obvious that a softer EOS might not be more susceptible to buoyant effects.
Acknowledgements.
The authors thank Elias Most, Patrick Cheong, and Eduardo M. Gutiérrez for helpful discussions. JF and DR acknowledge support from the U.S. Department of Energy, Office of Science, Division of Nuclear Physics under Award Numbers DE-SC0021177. DR also acknowledges funding from the National Science Foundation under Grants No. PHY-2020275, AST-2108467, PHY-2407681, the Sloan Foundation, and from the U.S. Department of Energy, Office of Science, Division of Nuclear Physics under Award Number DE-SC0024388. PH acknowledges funding from the National Science Foundation under Grant No. PHY-2116686. This work was completed in part at the TACC Open Hackathon, part of the Open Hackathons program. The authors would like to acknowledge OpenACC-Standard.org for their support. We are particularly grateful to Matthew Cawood, Victor Eijkhout, and Forrest Glines for their assistance optimizing AthenaK. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC0205CH11231.
Appendix A Equation of State Slices
For the hot remnant profile, we perform nonlinear fits to the temperature $T$ and lepton fraction $Y_{l}$ using the models
$T(x)$ $=A+Bx+CD\left[D+\left(\frac{x-x_{0}}{1-F\operatorname{sign}\left(x-x_{0}\right)}\right)^{2}\right]^{-1},$ (16a)
$Y_{l}(x)$ $=\sum_{i=0}^{7}c_{i}x^{i},$ (16b)
where $x=\log{\rho}$. We use 100 profiles of the $z$-axis equally spaced in time from the last $4.7$ ms of the remnant studied in Ref. [11] (${\sim}100~{}\mathrm{ms}$ in duration). For fits of $Y_{l}$, we compute $Y_{l}=Y_{e}+Y_{\nu_{e}}-Y_{\bar{\nu}_{e}}$, where $Y_{e}$, $Y_{\nu_{e}}$, and $Y_{\bar{\nu}_{e}}$ are the electron, electron neutrino, and electron antineutrino fractions, respectively. These fits are shown in the left half of Fig. 11.
This profile is extrapolated to high density based on these fits. This is necessary because the source profile is rotating, so the TOV case must be more compact to achieve a high-mass remnant (which is essential so that the shear layer is close enough to the surface to achieve buoyancy).
Figure 11: (Left) Fits of $T$ and $Y_{l}$ as functions of $\rho$ to the remnant from Ref. [11]. Black lines indicate profiles along the z-axis from the last $4.7$ msof the remnant, the red line marks the fit to this line. (Right) Mass-radius curves for the hot DD2 profile used in this work and the standard cold beta-equilibrium slice of DD2.
The mass-radius curve is shown in the right half of Fig. 11. For reference, we also plot the standard cold beta-equilibrated slice of DD2.
Appendix B Relativistic Magnetic Buoyancy in the TOV Spacetime
Note: for this section, we consider geometrized units with $G=c=k_{b}=1$ and adopt the Heaviside-Lorentz convention for the magnetic field; that is, $B^{i}_{\mathrm{HL}}=B^{i}_{\mathrm{CGS}}/\sqrt{4\pi}$. We also use the Einstein summation convention for paired indices, e.g., $a^{\mu}b_{\mu}\equiv\sum_{\mu}a^{\mu}b_{\mu}$.
Consider the relativistic generalization of Eq. 1 [44]:
$\left(e+P+b^{2}\right)\frac{du^{\alpha}}{d\tau}+\mathcal{P}^{\alpha}_{\nu}g^{\mu\nu}\partial_{\mu}\left(P+\frac{b^{2}}{2}\right)-\mathcal{P}^{\alpha}_{\nu}\nabla_{\mu}\left(b^{\mu}b^{\nu}\right)=-\left(e+P+b^{2}\right)\Gamma^{\alpha}_{\mu\nu}u^{\mu}u^{\nu},$ (17)
where $e$ is the total energy density, $P$ is the pressure, $b^{\mu}$ is the magnetic field as measured in the comoving frame, $u^{\alpha}$ is the four-velocity, $\tau$ is the proper time, and $\mathcal{P}^{\mu\nu}=g^{\mu\nu}+u^{\mu}u^{\nu}$ is a projection operator. Assume there exists a state in hydrostatic equilibrium with energy density $e_{0}$, pressure $P_{0}$, and no magnetic field. Then it follows that
$g^{\alpha i}\partial_{i}P_{0}=-\left(e_{0}+P_{0}\right)\Gamma^{\alpha}_{tt}u^{t}u^{t}.$ (18)
Now consider a state with $P^{\ast}=P_{0}-\Delta P$, $e^{\ast}=e_{0}-\Delta e$, and a magnetic field $b^{\mu}$. We again enforce $P_{0}=P^{\ast}+b^{2}/2$, which implies $\Delta P=b^{2}/2$. If we consider the initial moment when the fluid is still at rest ($u^{i}=0$, but derivatives of $u^{\mu}$ do not vanish), then Eq. 17 becomes
$\left(e^{\ast}+P^{\ast}+b^{2}\right)\frac{du^{\alpha}}{d\tau}-\nabla_{\mu}\left(b^{\mu}b^{\alpha}\right)-u^{\alpha}u_{\nu}b^{\mu}\nabla_{\mu}b^{\nu}=\left(\Delta e-\frac{b^{2}}{2}\right)\Gamma^{\alpha}_{tt}u^{t}u^{t}.$ (19)
We now focus exclusively on $u^{r}$, which eliminates the third term on the left-hand side. Expanding the covariant derivative leads to
$\left(e^{\ast}+P^{\ast}+b^{2}\right)\frac{du^{r}}{d\tau}-\partial_{\mu}\left(b^{\mu}b^{r}\right)-\Gamma^{\mu}_{\mu\nu}b^{\nu}b^{r}-\Gamma^{r}_{\mu\nu}b^{\mu}b^{\nu}=\left(\Delta e-\frac{b^{2}}{2}\right)\Gamma^{r}_{tt}u^{t}u^{t}.$ (20)
As in the Newtonian case, we assume only the azimuthal field, $b^{\phi}$, is nonzero (note that the $b^{t}$ term vanishes if $u_{i}=0$). This causes the $\partial_{\mu}\left(b^{\mu}b^{r}\right)$ term to vanish, leading to
$\left(e^{\ast}+P^{\ast}+b^{2}\right)\frac{du^{r}}{d\tau}-\Gamma^{r}_{\phi\phi}b^{\phi}b^{\phi}=\left(\Delta e-\frac{b^{2}}{2}\right)\Gamma^{r}_{tt}u^{t}u^{t}.$ (21)
In the TOV spacetime, $\Gamma^{r}_{\phi\phi}=-\left(1-2m(r)/r\right)r\sin^{2}\theta$ and $\Gamma^{r}_{tt}=\left(1-2m(r)/r\right)\alpha\partial_{r}\alpha$, where $\alpha$ is the lapse function. Since $b^{\phi}$ is the only nonzero magnetic field component, $b^{2}=\left(b^{\phi}\right)^{2}r^{2}\sin^{2}\theta$. Lastly, $u^{0}u^{0}g_{00}=-1$ implies that $u^{0}=1/\sqrt{-g_{00}}$, so $u^{0}=1/\alpha$. Therefore,
$\frac{b^{2}}{r}<\left(\Delta e-\frac{b^{2}}{2}\right)\frac{\partial_{r}\alpha}{\alpha}.$ (22)
From the TOV equations,
$\frac{d\alpha}{dr}=\frac{m(r)+4\pi r^{3}P_{0}}{r^{2}\left(1-2m(r)/r\right)}\alpha$ (23)
so
$b^{2}<\left(\Delta e-\frac{b^{2}}{2}\right)\frac{m(r)+4\pi r^{3}P_{0}}{r-2m(r)}$ (24)
We will now show that this reduces to the Newtonian limit. First, recall that $b^{2}/2=\Delta P$ and $e=\rho+\rho\epsilon$. Then $\Delta e-b^{2}/2=\Delta\rho+\Delta(\rho\epsilon)-\Delta P$. Adding back in the missing factors of $c^{2}$, the right-hand side of Eq. 24 becomes
$\left(\Delta\rho+\frac{1}{c^{2}}\Delta(\rho\epsilon)-\frac{\Delta P}{c^{2}}\right)\frac{m(r)}{r}\frac{\left(1+4\pi r^{3}\frac{P_{0}}{m(r)c^{2}}\right)}{\left(1-\frac{2m(r)}{rc^{2}}\right)}$ (25)
In the non-relativistic limit, $c\rightarrow\infty$ and $b^{i}\rightarrow B^{i}$, leaving us with
$B^{2}<\Delta\rho\frac{m(r)}{r}.$ (26)
Returning to CGS units, this yields
$B^{2}<4\pi\Delta\rho\frac{Gm(r)}{r},$ (27)
which agrees with Eq. 5.
Eq. 24 is often a weaker constraint than its Newtonian counterpart. A massive remnant may have $m(r)\sim 2~{}\textrm{M}_{\odot}$ near a shear layer located at $r\sim 2~{}\textrm{M}_{\odot}$, suggesting that $(m(r)+4\pi r^{3}P_{0})/(r-2m(r))$ can be more than a factor of $2$ larger than $m(r)/r$. The precise behavior of the term $\Delta e-\Delta P$ depends on the actual EOS, but a sufficient condition to ensure that $\Delta e-\Delta P>\Delta\rho$ is that $d(\rho\epsilon)/dP>1$. This is always satisfied for an ideal gas respecting causality ($1<\Gamma\leq 2$), though it may not always hold for a realistic EOS. Nevertheless, we have found for our cases that the product $(\Delta e-\Delta P)(1+r\pi r^{3}P_{0}/m(r))(1-2m(r)/r)^{-1}>1$.

### v2

@@
 I Introduction
-During a binary neutron star (BNS) merger, the shearing motion between the stellar surfaces as they contact triggers turbulent instabilities, which help amplify the magnetic field to magnetar strengths within the first few milliseconds after merger [1, 2, 3, 4, 5, 6].
-In the post-merger phase, this shear layer where the KHI is active settles to a region ${\sim}1~{}\mathrm{km}$ below the surface [7, 8], which is where the strongest fields are focused [5, 6].
-Recent simulations suggest these strongly amplified fields can break out from the remnant star on relatively short timescales and help power short gamma-ray bursts (sGRBs) [9, 10].
-However, it is not immediately clear how this breakout occurs; BNS remnants are typically stable against convection [7, 8, 11, 12], so this breakout cannot be purely hydrodynamical in nature.
+During a binary neutron star (BNS) merger, the shearing motion between the stellar surfaces as they contact triggers turbulent instabilities, which help amplify the magnetic field to magnetar strengths within the first few milliseconds after merger Price and Rosswog (2006); Kiuchi et al.
+(2014, 2015, 2018); Palenzuela et al.
+(2022); Aguilera-Miret et al.
+(2023).
+In the post-merger phase, this shear layer where the Kelvin-Helmholtz instability is active settles to a region ${\sim}1~\mathrm{km}$ below the surface Kastaun and Galeazzi (2015); Hanauske et al.
+(2017), which is where the strongest fields are focused Palenzuela et al.
+(2022); Aguilera-Miret et al.
+(2023).
+Recent simulations suggest these strongly amplified fields can break out from the remnant star on relatively short timescales and help power short gamma-ray bursts (sGRBs) Most and Quataert (2023); Combi and Siegel (2023).
+However, it is not immediately clear how this breakout occurs; BNS remnants are typically stable against convection Kastaun and Galeazzi (2015); Hanauske et al.
+(2017); Radice and Bernuzzi (2023); Gao et al.
+(2026), so this breakout cannot be purely hydrodynamical in nature.
 A similar problem occurs in solar physics.
@@
 Their progress then stalls inside the photosphere, which is stable against convection.
-A rather successful model for achieving emergence relies on magnetic buoyancy: as magnetic flux accumulates near the base of the photosphere, the magnetized region eventually becomes buoyant and accelerates upward [13].
+A rather successful model for achieving emergence relies on magnetic buoyancy: as magnetic flux accumulates near the base of the photosphere, the magnetized region eventually becomes buoyant and accelerates upward Archontis et al.
+(2004).
 To illustrate how this occurs, consider a pocket of fluid with a magnetic field $\mathbf{B}$ in pressure equilibrium with an ambient unmagnetized fluid.
@@
 A strong toroidal magnetic field develops inside the shear layer (see, e.g., Refs.
-[14, 15]), which could become buoyant as it settles into hydrostatic equilibrium with the surrounding fluid.
+Aguilera-Miret et al.
+(2022); Kiuchi et al.
+(2024)), which could become buoyant as it settles into hydrostatic equilibrium with the surrounding fluid.
 From the Newtonian ideal magnetohydrodynamics (MHD) equations, the velocity $\mathbf{v}$ of the flux tube evolves as
@@
 If we assume there exists a magnetic flux tube with $\mathbf{B}\neq 0$, pressure $P^{\ast}$ satisfying $P_{0}=P^{\ast}+B^{2}/8\pi$ (i.e., pressure equilibrium), and density $\rho^{\ast}$, the velocity is described by
-$\rho^{\ast}\frac{D\mathbf{v}}{Dt}=\left(\mathbf{B}\cdot\nabla\right)\mathbf{B}-\Delta\rho\mathbf{g},$ (2)
+$\rho^{\ast}\frac{D\mathbf{v}}{Dt}=\frac{1}{4\pi}\left(\mathbf{B}\cdot\nabla\right)\mathbf{B}-\Delta\rho\mathbf{g},$ (2)
 where $\Delta\rho=\rho_{0}-\rho^{\ast}$. For a spherically symmetric neutron star, we can specialize to spherical coordinates and the $r$-direction. Therefore,
@@
 This differs slightly from buoyancy in solar physics, which typically concerns relatively small sections of magnetized flux tubes.
-In the solar regime, a parallel-plane approximation is appropriate (e.g., [16, 17, 18, 13, 19, 20]), so there is no tension force at all.
+In the solar regime, a parallel-plane approximation is appropriate (e.g., Matsumoto et al.
+(1993); Longcope et al.
+(1996); Magara and Longcope (2001); Archontis et al.
+(2004); Murray et al.
+(2006); Toriumi et al.
+(2011)), so there is no tension force at all.
 In the case of a BNS remnant, however, we are interested in the large-scale circumstellar toroidal field, and tension cannot be neglected.
@@
 However, a semi-relativistic treatment (i.e., strong gravity and large internal energy but $v/c\ll 1$, see Appendix B) leads to a modified buoyancy condition which is usually less restrictive.
-Note that this condition only determines buoyancy at a given moment and does not necessarily mean a toroidal field inside a BNS remnant will emerge; similar to classic hydrodynamic buoyancy, a stably stratified atmosphere can impede magnetic buoyancy, and finite-size effects (e.g., drag) as the scale height shrinks can also affect buoyancy [21, 16].
+Note that this condition only determines buoyancy at a given moment and does not necessarily mean a toroidal field inside a BNS remnant will emerge; similar to classic hydrodynamic buoyancy, a stably stratified atmosphere can impede magnetic buoyancy, and finite-size effects (e.g., drag) as the scale height shrinks can also affect buoyancy Spruit et al.
+(1987); Matsumoto et al.
+(1993).
 Nevertheless, this condition suggests that magnetic buoyancy may be a possible mechanism for embedded magnetic fields in BNS remnants to emerge.
@@
 V.
+Though most quantities in this paper are expressed in physical units, a number of formulae in Sec.
+III.1 and App.
+B instead use geometrized units with $G=c=1$ to maintain consistency with relativistic literature.
+In these scenarios, the solar mass $M_{\odot}$ is chosen as the relevant dimensional scale.
+However, any results obtained with these formulae quoted in the text are first converted to physical units.
 II Methods
@@
 We construct our model based on the results of Ref.
-[11].
+Radice and Bernuzzi (2023).
 As shown in Figures 3, 7, and 8 of Ref.
-[11], the remnant settles into a quasi-steady state approximately $10~{}\mathrm{ms}$ after merger.
-We perform nonlinear fits as functions of density (see Appendix A) to the temperature and lepton fraction profiles in order to compute a 1D finite-temperature equilibrium slice of the HS(DD2) EOS (hereafter DD2)[22, 23], which we then use to generate a static neutron star via the Tolman-Oppenheimer-Volkoff equations [24, 25].
-Using a central density $\rho=2.2\times 10^{-3}~{}\textrm{M}_{\odot}^{-2}\approx 1.36\times 10^{15}~{}\mathrm{g}/\mathrm{cm}^{3}$, the resulting star has a mass of $M\approx 2.43~{}\textrm{M}_{\odot}$ and a radius $R\approx 8.35~{}\textrm{M}_{\odot}=12.3~{}\mathrm{km}$.
-Because our table only covers a finite density range, we define the edge of the star as the point dropping below $\rho=10^{-5}~{}\textrm{M}_{\odot}^{-2}\approx 6.18\times 10^{13}~{}\mathrm{g}/\mathrm{cm}^{3}$.
-Because we are interested in the dynamics of buoyant magnetic structures, we add a toroidal flux tube of radius $s_{0}$ satisfying
+Radice and Bernuzzi (2023), the remnant settles into a quasi-steady state approximately $10~\mathrm{ms}$ after merger.
+We perform nonlinear fits as functions of density (see Appendix A) to the temperature and lepton fraction profiles in order to compute a 1D finite-temperature equilibrium slice of the HS(DD2) EOS (hereafter DD2)Typel et al.
+(2010); Hempel and Schaffner-Bielich (2010), which we then use to generate a static neutron star via the Tolman-Oppenheimer-Volkoff (TOV) equations Tolman (1939); Oppenheimer and Volkoff (1939).
+Using a central density $\rho=2.2\times 10^{-3}~G^{-3}\textrm{M}_{\odot}^{-2}c^{6}\approx 1.36\times 10^{15}~\mathrm{g}/\mathrm{cm}^{3}$, the resulting star has a mass of $M\approx 2.43~\textrm{M}_{\odot}$ and a radius $R\approx 8.35~G\textrm{M}_{\odot}c^{-2}\approx 12.3~\mathrm{km}$.
+Because our table only covers a finite density range, we define the edge of the star as the point dropping below $\rho=10^{-5}~G^{-3}\textrm{M}_{\odot}^{-2}c^{6}\approx 6.18\times 10^{12}~\mathrm{g}/\mathrm{cm}^{3}$.
+Because we are interested in the dynamics of buoyant magnetic structures, we modify the initial data, which is initially in hydrostatic equilibrium, by superimposing a toroidal flux tube of radius $s_{0}$ satisfying
 $\sqrt{\gamma}B^{\phi}$ $=B_{0}e^{-s^{2}/\sigma^{2}}$ (8a)
@@
 A realistic magnetic field in a merger is not completely axisymmetric, which can trigger additional magnetic buoyant instabilities beyond the simple form of magnetic buoyancy already discussed.
-Among the most important of these effects is the so-called Parker instability (or undular instability) [26, 27].
+Among the most important of these effects is the so-called Parker instability (or undular instability) Parker (1966); Acheson (1979).
 Consider a longitudinal perturbation (i.e., parallel to the field lines) to the density or velocity to a state otherwise in equilibrium.
@@
 along the length of the tube where $A$ is a specified amplitude and $k$ is an integer determining the harmonic; because we keep the pressure fixed, this indirectly adds a perturbation to the temperature inside the flux tube.
-These longitudinal perturbations add a small amount of hydrodynamic buoyancy to the system, causing the field lines to bend and possibly triggering the Parker instability.
-After constructing the flux tube, we enforce mechanical equilibrium by setting the tube’s pressure to $P^{\ast}=P_{0}-P_{B}$, where $P_{0}$ is the ambient fluid pressure and $P_{B}$ is magnetic pressure ($P_{B}=b^{2}/8\pi$ in CGS units, where $b^{\mu}$ is the magnetic field in the fluid’s rest frame).
-We then reset the fluid density to enforce thermal equilibrium with the (perturbed) ambient medium to create a buoyant force.
+These longitudinal perturbations add a small amount of hydrodynamic buoyancy to the system, causing the field lines to bend and possibly seeding the Parker instability.
+After superimposing the magnetic flux tube and applying the fluid perturbation, we enforce mechanical equilibrium with the TOV solution by setting the pressure inside the tube to $P^{\ast}=P_{0}-P_{B}$, where $P_{0}$ is the ambient fluid pressure and $P_{B}$ is magnetic pressure ($P_{B}=b^{2}/8\pi$ in CGS units, where $b^{\mu}$ is the magnetic field in the fluid’s rest frame).
+The density, which has already been perturbed, is then adjusted again to keep the (perturbed) temperature fixed.
+The composition of the fluid is also held constant during this process.
+Therefore, the fluid is in pressure equilibrium with the TOV solution and thermal equilibrium with the perturbed state, but it deviates slightly from thermal equilibrium with the TOV (leading to hydrodynamic buoyancy), and it cannot generally achieve hydrostatic equilibrium due to the magnetic field.
 In a true post-merger scenario, the thermal diffusion time scale is very long compared to the dynamical time scale, so calculating the density difference assuming an adiabatic process rather than an isothermal one may produce a more realistic flux tube.
@@
 Table 1: A summary of the simulations performed, including the initial maximum field strength, the radius $s_{0}$ of the flux tube, the estimated emergence time, and the estimated velocity at emergence.
-Name $\sqrt{b^{2}_{\mathrm{max}}}$ (G) $A$ $s_{0}$ ($\textrm{M}_{\odot}$) $t_{\mathrm{em}}$ (ms) $v_{\mathrm{em}}$
-H7e17_5 $7.540\times 10^{17}$ $0.05$ $0.25$ 0.216 0.038
-H7e17_5_B $7.540\times 10^{17}$ $0.05$ $1.0$ 0.201 0.013
-H7e17_0 $7.540\times 10^{17}$ 00 $0.25$ 0.224 0.034
-H1e17_5 $1.257\times 10^{17}$ $0.05$ $0.25$ N/A N/A
-H1e17_0 $1.257\times 10^{17}$ 00 $0.25$ N/A N/A
-H1e16_5 $1.257\times 10^{16}$ $0.05$ $0.25$ N/A N/A
-H6e15_5 $6.283\times 10^{15}$ $0.05$ $0.25$ N/A N/A
-H0_5 00 $0.05$ $0.25$ N/A N/A
-For most of our runs, we consider a flux tube of inner radius $s_{0}=0.25~{}\textrm{M}_{\odot}\approx 369~{}\mathrm{m}$ and an outer radius $R$ matching the position of the peak temperature ($r=7.114~{}\textrm{M}_{\odot}$), which we take as a proxy for the position of the shear layer.
-We additionally set the width parameter $\sigma=0.3$ and add a small twist with $q=1$ to increase rigidity of the flux tube.
+Name $\sqrt{b^{2}_{\mathrm{max}}}$ (G) AA $s_{0}$ (km) k $t_{\mathrm{em}}$ (ms) $v_{\mathrm{em}}(c)$
+H7e17_5 $7.540\times 10^{17}$ $0.05$ $0.369$ 18 0.216 0.038
+H7e17_5_B $7.540\times 10^{17}$ $0.05$ $1.48$ 18 0.201 0.013
+H7e17_5_k9 $7.540\times 10^{17}$ $0.05$ $0.369$ 9 0.194 0.048
+H7e17_0 $7.540\times 10^{17}$ 0 $0.369$ N/A 0.224 0.034
+H1e17_5 $1.257\times 10^{17}$ $0.05$ $0.369$ 18 N/A N/A
+H1e17_5_k9 $1.257\times 10^{17}$ $0.05$ $0.369$ 9 N/A N/A
+H1e17_0 $1.257\times 10^{17}$ 0 $0.369$ N/A N/A N/A
+H1e16_5 $1.257\times 10^{16}$ $0.05$ $0.369$ 18 N/A N/A
+H6e15_5 $6.283\times 10^{15}$ $0.05$ $0.369$ 18 N/A N/A
+H0_5 0 $0.05$ $0.369$ 18 N/A N/A
+For most of our runs, we consider a flux tube of inner radius $s_{0}=0.25~G\textrm{M}_{\odot}c^{-2}\approx 369~\mathrm{m}$ and an outer radius $R$ matching the position of the peak temperature ($r=7.114~\textrm{M}_{\odot}$), which we take as a proxy for the position of the shear layer.
+We additionally set the width parameter $\sigma=0.3$ and add a small twist with $q=1~G^{-1}\textrm{M}_{\odot}^{-1}c^{2}\approx 0.677~\mathrm{km}^{-1}$ to increase rigidity of the flux tube.
 The perturbation wavenumber is set to $k=18$.
@@
 As a control, we also perform a run with no magnetic field but with an $A=0.05$ density perturbation; this will help disentangle the effects of ordinary hydrodynamic buoyancy from those caused by the Parker instability.
-To help track the evolution of the tube for this case, a passive tracer scalar $s$ is added with $s=1$ inside the tube and $s=0$ everywhere else.
+To help track the evolution of the tube for this case, a passive tracer scalar $X$ is added with $X=1$ inside the tube and $X=0$ everywhere else.
 Each run is recorded in Table 1, with the names following the format “Px_A”.
@@
 Thus the run with a $7.540\times 10^{17}~\textrm{G}$ field embedded in the hot remnant profile with an $A=0.05$ perturbation is labeled “H7e17_5“, while “H0_5” corresponds to the base case with a density perturbation but no magnetic field.
-Since the shear layer is ${\sim}3~{}\mathrm{km}$ wide in a realistic BNS merger [8], a flux tube of $s_{0}=0.25~{}\textrm{M}_{\odot}$ may be much smaller in size than the toroidal field structure which develops in a realistic merger simulation.
+Since the shear layer is ${\sim}3~\mathrm{km}$ wide in a realistic BNS merger Hanauske et al.
+(2017), a flux tube of $s_{0}=0.25~G\textrm{M}_{\odot}/c^{2}$ may be much smaller in size than the toroidal field structure which develops in a realistic merger simulation.
 Because we may approximate the field as a superposition of weakly twisted flux tubes, tests with small flux tubes should be qualitatively correct when determining if a particular flux tube can emerge.
 Nevertheless, we do include one additional setup with the H7e17_5 configuration with a larger flux tube as well.
-For this run, we set $s_{0}=1.0~{}\textrm{M}_{\odot}\approx 1.48~{}\mathrm{km}$ and reduce the twist coefficient to $q=0.25$; since the twist scales linearly with the tube radius $s$, this is necessary to ensure the magnetic field remains primarily toroidal.
+For this run, we set $s_{0}=1.0~G\textrm{M}_{\odot}/c^{2}\approx 1.48~\mathrm{km}$ and reduce the twist parameter to $q=0.25~G^{-1}\textrm{M}_{\odot}^{-1}c^{2}\approx 0.169~\mathrm{km}^{-1}$; since the twist scales linearly with the tube radius $s$, this is necessary to ensure the magnetic field remains primarily toroidal.
 We label this run H7e17_5_B to distinguish it from the case with a small flux tube.
-We note that our strongest fields are unrealistically large for a BNS post-merger; high-resolution BNS simulations starting with realistic initial field strengths do not typically produce fields with peak strengths much in excess of $10^{17}~{}\textrm{G}$, and average strengths are typically ${\sim}10^{16}~{}\textrm{G}$ [3, 4, 28, 5, 14, 15].
-Though no full-scale merger simulation has sufficient resolution to capture all scales of magnetic turbulence, local simulations confirm these findings [29, 30].
+We also consider two additional runs similar to H7e17_5 and H7e17_1 but with $k=9$, which we label as H7e17_5_k9 and H1e17_5_k9, respectively, to better understand any possible role of the Parker instability, which is wavenumber-dependent.
+In a plane-parallel approximation in Newtonian MHD, the maximum growth rate occurs at $\lambda\sim 10H_{P}$, where $\lambda$ is the wavelength and $H_{P}$ is the pressure scale height Horiuchi et al.
+(1988).
+This corresponds to $k\approx 8$ for our configuration, so we choose $k=9$ as it is the smallest perturbation that fits a full wavelength inside our domain.
+We acknowledge that our system is neither plane-parallel nor Newtonian, but we expect that the leading-order corrections to the growth rate from GR and a toroidal geometry should be similar across all wavenumbers.
+Note also that $k\approx 8$ is only the critical wavenumber near the initial location of the flux tube.
+As the flux tube rises, the scale height will rapidly decrease, so the critical wavenumber will increase.
+The “correct” perturbation for maximum growth, then, depends on the timescale for the Parker instability during the linear phase.
+If this growth rate is much shorter than the emergence timescale, the linear phase will be dominated by the initial conditions.
+At the other extreme, where the timescale is much longer than the emergence timescale, the Parker instability will not play a significant role in the dynamics, so the perturbation will be irrelevant.
+We note that our strongest fields are unrealistically large for a BNS post-merger; high-resolution BNS simulations starting with realistic initial field strengths do not typically produce fields with peak strengths much in excess of $10^{17}~\textrm{G}$, and average strengths are typically ${\sim}10^{16}~\textrm{G}$ Kiuchi et al.
+(2015, 2018); Mösta et al.
+(2020); Palenzuela et al.
+(2022); Aguilera-Miret et al.
+(2022); Kiuchi et al.
+(2024).
+Though no full-scale merger simulation has sufficient resolution to capture all scales of magnetic turbulence, local simulations confirm these findings Obergaulinger et al.
+(2010); Zrake and MacFadyen (2013).
+Note also that the remnant in Ref.
+Radice and Bernuzzi (2023) is differentially rotating, but we have neglected these effects by using a TOV solution.
+Though rotation can modify buoyancy and convective instabilities, it tends to inhibit these mechanisms Acheson and Gibbons (1978); Spruit (1999); Gao et al.
+(2026).
+In conjunction with our assumption of a flux tube in thermal (rather than isentropic) equilibrium, our results will tend toward overestimating the buoyancy of the system.
 II.2 Numerical Methods
-We evolve the star with the GRMHD code AthenaK [31, 32, 33], which uses a second-order finite-volume method to evolve the fluid and upwind constrained transport to ensure that $\nabla\cdot\mathbf{B}=0$.
-For these tests, we choose WENOZ as our reconstruction scheme [34] and enable a first-order flux correction (FOFC) method [35].
+We evolve the star with the GRMHD code AthenaK Stone et al.
+(2024); Zhu et al.
+(2025); Fields et al.
+(2025), which uses a second-order finite-volume method to evolve the fluid and upwind constrained transport to ensure that $\nabla\cdot\mathbf{B}=0$.
+For these tests, we choose WENOZ as our reconstruction scheme Borges et al.
+(2008) and enable a first-order flux correction (FOFC) method Lemaster and Stone (2009).
 The FOFC scheme greatly improves the stability of the evolution near the surface of the neutron star, particularly when the magnetization is high.
 Because we assume that the remnant has achieved steady state, we fix the spacetime to the initial TOV solution.
-Our computational domain is a wedge in the equatorial plane covering $r\in[6~{}\textrm{M}_{\odot},24~{}\textrm{M}_{\odot}]$, $\theta\in[\pi/2-\pi/9,\pi/2+\pi/9]$, and $\phi\in[-\pi/9,\pi/9]$, which amounts to full two oscillations for the $k=18$ perturbation described in Sec.
+Because the energy density inside the flux tube is reduced, there will be a gravitational mass deficit that should affect the metric.
+However, even for the H7e17_5_B case, this deficit is $\lesssim 2.2\%$ (and $\lesssim 0.2\%$ for H7e17_5), so this effect is a small perturbation that can safely be ignored.
+Our computational domain is a wedge in the equatorial plane covering $r\in[6~G\textrm{M}_{\odot}c^{-2},24~G\textrm{M}_{\odot}c^{-2}]\approx[8.86~\mathrm{km},35.44~\mathrm{km}]$, $\theta\in[\pi/2-\pi/9,\pi/2+\pi/9]$, and $\phi\in[-\pi/9,\pi/9]$, which amounts to full two oscillations for the $k=18$ perturbation described in Sec.
 II.1.
-We use a base domain of $600\times 180\times 180$ cells with a single refinement region over $r\in[6~{}\textrm{M}_{\odot},12\textrm{M}_{\odot}]$ for an effective radial resolution of $\Delta r=0.015~{}\textrm{M}_{\odot}\approx 22~{}\mathrm{m}$ inside the star.
-The angular resolution is sufficient that $r\Delta\theta\approx r\sin{\theta}\Delta\phi\approx\Delta r$ over the region of interest from approximately $7~{}\textrm{M}_{\odot}$ to $9~{}\textrm{M}_{\odot}$.
-The upper boundary of $r$ is a “diode” outflow boundary (the radial momentum must be nonnegative), and the $\theta$ and $\phi$ boundaries are periodic.
-For the inner boundary of $r$, we keep the rest-mass density $\rho$ and pressure $P$ fixed, but we apply a reflecting boundary to the magnetic field $B^{i}$ and the radial fluid momentum density $\alpha\sqrt{\gamma}\rho Wv^{r}$, where $\gamma$ is the determinant of the spatial metric, $\alpha$ is the lapse, $v^{r}$ is the radial component of the fluid three-velocity, and $W=(1-v^{2})^{-1/2}$ is the Lorentz factor.
-This condition ensures that the density flux across the boundary is approximately zero.
-The $Wv^{\phi}$ and $Wv^{\theta}$ velocity components are also reflected across the boundary with the parity inverted.
-Though these components are parallel to the boundary, we found that inverting the parity improves stability; otherwise, small numerical errors can accumulate in the velocity, and a simple reflection can amplify these errors in a way that gives the star an unphysical rotation, while flipping the sign of the velocity components damps them.
-This work represents the first use of 3D EOS tables in AthenaK, which are implemented following the same prescription as in GR-Athena++ [36].
+We use a base domain of $600\times 180\times 180$ cells with a single refinement region over $r\in[6~G\textrm{M}_{\odot}c^{-2},12~G\textrm{M}_{\odot}c^{-2}]\approx[8.86~\mathrm{km},17.72~\mathrm{km}]$ for an effective radial resolution of $\Delta r=0.015~G\textrm{M}_{\odot}c^{-2}\approx 22~\mathrm{m}$ inside the star.
+The angular resolution is sufficient that $r\Delta\theta\approx r\sin{\theta}\Delta\phi\approx\Delta r$ over the region of interest from approximately $7~G\textrm{M}_{\odot}c^{-2}\approx 10~\mathrm{km}$ to $9~G\textrm{M}_{\odot}c^{-2}\approx 13~\mathrm{km}$.
+The upper boundary of $r$ is a “diode” outflow boundary (the radial momentum must be nonnegative), which limits spurious fluxes generated at the level of the artificial atmosphere from inflowing back into the computational domain.
+The inner boundary of $r$ is designed to approximate a “stress-free” boundary.
+We first fix the rest-mass density $\rho$ and pressure $P$ to the initial TOV solution.
+The radial velocity $Wv^{r}$, where $W=(1-v^{2})^{-1/2}$ is the Lorentz factor and $v^{i}$ is the three-velocity, is then determined by setting $F_{\mathrm{in}}=-F_{\mathrm{out}}$ for the mass flux $F=\alpha\sqrt{\gamma}\rho Wv^{r}$, where $\gamma$ is the determinant of the spatial metric and $\alpha$ is the lapse, and solving for $(Wv^{r})_{\mathrm{in}}$.
+This behaves nearly like a typical reflection boundary, but it corrects for $\alpha\sqrt{\gamma}\rho$ being slightly larger on the inside of the star.
+The magnetic field and $Wv^{\phi}$ and $Wv^{\theta}$ are also reflected.
+Note, however, that we invert the parity for all components rather than only the radial component.
+In the event that the flux tube comes too close to the inner boundary, we found that a simple reflection that does not invert the parity of the tangential components of the magnetic field and velocity can cause small numerical errors to accumulate in the velocity and generate an unphysical rotation in the star.
+Through experimentation, we found that also flipping the tangential components damps these errors out and greatly improves stability.
+In practice the flux tube remains far from the radial boundary in all the runs presented here, so this condition has little effect on the evolution.
+We exploit the spherical symmetry of the TOV solution and apply periodic boundaries in both the $\theta$ and $\phi$ directions.
+Though this means there are multiple copies of the flux tube along the polar direction, this is only an issue if the flux tube expands enough to interact with the polar boundary.
+So long as the flux tube is sufficiently resolved, this only happens some time after the flux tube emerges from the star; we end our runs before this occurs.
+This work represents the first use of 3D EOS tables in AthenaK, which are implemented following the same prescription as in GR-Athena++ Cook et al.
+(2025a).
 To approximate their effects on the remnant, we modify the full DD2 EOS table with the contributions of trapped neutrinos following the prescription in Ref.
-[37].
+Perego et al.
+(2019).
 Therefore, the evolved species fraction is now the (electron) lepton fraction, $Y_{l}=Y_{e}+(n_{\nu_{e}}-n_{\nu_{\overline{e}}})/n_{b}$, which is advected along with the fluid flow.
@@
 However, because the remnant is stable against convection, the perturbations slow down and eventually settle into layers of the remnant with similar densities by $t=0.39~\textrm{ms}$.
-Figure 1: A slice of the $r$-$\phi$ plane at $t=0.39~{}\textrm{ms}$ comparing $|B|$ from H1e16_5 (left) to the tracer scalar $s_{1}$ from H0_5 (right).
+Figure 1: A slice of the $r$-$\phi$ plane at $t=0.39~\textrm{ms}$ comparing $|B|$ from H1e16_5 (left) to the tracer scalar $X$ from H0_5 (right).
 The dashed black line marks the surface of the star, and the dot-dashed green lines mark the boundaries of the initial position of the flux tube.
@@
 Similar to Ref.
-[38], we express the kinetic energy $E_{K}$ and magnetic energy $E_{B}$ as
+Cook et al.
+(2025b), we express the kinetic energy $E_{K}$ and magnetic energy $E_{B}$ as
 $E_{K}$ $=\int\rho W\left(W-1\right)\mathrm{d}V,$ (13)
@@
 Since $\tau$ does not include the gravitational energy of the system, it is not conserved.
-However, $-\alpha(T^{0}_{0}+\rho u^{0})$ does include these contributions and is therefore conserved in stationary spacetimes.
+However, $\alpha(T^{0}_{0}+\rho u^{0})$ does include these contributions and is therefore conserved in stationary spacetimes.
 It is easy to verify that $\alpha T^{0}_{0}+\alpha\rho u^{0}+\tau\approx-\rho\Phi$ in the Newtonian limit, where $\Phi$ is the gravitational potential.
@@
 which we use to normalize the energy of our system.
-We caution that $E_{\mathrm{tot}}$ is not precisely invariant in due to boundary effects and truncation error from evolving $\tau$ instead of $\alpha(T^{0}_{0}+\rho u^{0})$.
-However, we have confirmed that our results are the same if we normalize by $E(t=0)$ or $E(t)$, indicating that conservation violations are insignificant for our analysis.
+We caution that $E_{\mathrm{tot}}$ is not precisely invariant due to boundary effects and truncation error from evolving $\tau$ instead of $\alpha(T^{0}_{0}+\rho u^{0})$, leading to a maximum deviation of ${\sim}0.5\%$ in the total energy over the course of the evolution.
+However, we have confirmed that our results are the same if we normalize by $E(t=0)$ or $E(t)$, indicating that these conservation violations are insignificant for our analysis.
 Figure 2: The magnetic energy $E_{B}$ (solid lines) and kinetic energy $E_{K}$ (dashed lines) as a function of time normalized by the total energy for all the runs with perturbations.
@@
 The magnetic energy decays as well, albeit more slowly than the kinetic energy.
-If we consider a volume $V=A\ell$, where $A$ is some changing cross-sectional area and $\ell$ is an approximately fixed length, flux conservation implies that $E_{B}\approx B^{2}V\propto A^{-1}$, so we attribute this slow decay in magnetic energy to the slow diffusion or expansion of the flux tube as demonstrated in Fig.
+If we consider a volume $V=A\ell$, where $A$ is some changing cross-sectional area and $\ell$ is an approximately fixed length, flux conservation implies that $2E_{B}\approx B^{2}V\propto A^{-1}$, so we attribute this slow decay in magnetic energy to the slow diffusion or expansion of the flux tube as demonstrated in Fig.
 3.
@@
 4 of Ref.
-[39]).
+Toriumi and Yokoyama (2011)).
 As it reaches the surface, the flux tube rapidly expands and pushes out a wave of material in front of it (see Fig.
@@
 H7e17_5 expands rapidly as it emerges, while H1e17_5 remains trapped beneath the surface and largely intact.
-Figure 6: A comparison of $|B|$ (top) and $\rho$ (bottom) in the $r$-$\phi$ plane from H7e17_5 at $t=0.34~{}\textrm{ms}$.
+Figure 6: A comparison of $|B|$ (left) and $\rho$ (right) in the $r$-$\phi$ plane from H7e17_5 at $t=0.34~\textrm{ms}$.
 Much of the flux tube remains near the surface, but the wave of material pushed out by it is also visible.
 The dashed black line marks the location of the stellar surface from the TOV solution, and the dot-dashed green lines mark the boundaries of the flux tube at $t=0$.
-To measure the emergence time and properties of the entrained material, we integrate the mass flux $\dot{M}$ across surfaces at $r=8.35~{}\textrm{M}_{\odot}$, $r=8.4~{}\textrm{M}_{\odot}$, $r=8.5~{}\textrm{M}_{\odot}$, $r=8.6~{}\textrm{M}_{\odot}$, and $r=8.7~{}\textrm{M}_{\odot}$.
-We define the emergence time as the peak mass flux $\dot{M}$ across the surface at $r=8.35~{}\textrm{M}_{\odot}$, which is the approximate location of the remnant’s surface.
+To measure the emergence time and properties of the entrained material, we integrate the mass flux $\dot{M}$ across surfaces at $r=8.35~G\textrm{M}_{\odot}c^{-2}\approx 12.33~\mathrm{km}$, $r=8.4~G\textrm{M}_{\odot}c^{-2}\approx 12.40~\mathrm{km}$, $r=8.5~G\textrm{M}_{\odot}c^{-2}\approx 12.55~\mathrm{km}$, $r=8.6~G\textrm{M}_{\odot}c^{-2}\approx 12.70~\mathrm{km}$, and $r=8.7~G\textrm{M}_{\odot}c^{-2}\approx 12.85~\mathrm{km}$.
+We define the emergence time as the peak mass flux $\dot{M}$ across the surface at $r=8.35~G\textrm{M}_{\odot}c^{-2}$, which is the approximate location of the remnant’s surface.
 By looking for where $\dot{M}$ peaks in the other integration surfaces, we can measure the approximate speed of the flux tube at emergence, which we do by fitting a quadratic polynomial for $r(t)$ to these peak times and taking its time derivative.
@@
 The mass across each integration surface also rapidly decreases with distance, suggesting that little to no mass is unbound [^1] A better way to estimate unbound ejecta is using the geodesic and/or Bernoulli criteria, but this must be done farther away from the remnant to be accurate.
-Because our simulations only span $[-\pi/9,\pi/9]$ in the $\theta$ direction with periodic boundary conditions, the emerging fluid begins to interact with itself not long after passing the surface at $r=8.7~{}\textrm{M}_{\odot}$.
+Because our simulations only span $\pi/9,\pi/9]$ in the $\theta$ direction with periodic boundary conditions, the emerging fluid begins to interact with itself not long after passing the surface at $r=8.7~G\textrm{M}_{\odot}c^{-2}$.
 Therefore, the dynamics of the fluid will not be reliable at a radius where a proper ejecta calculation would be meaningful..
@@
 Unsurprisingly, Fig.
-2 demonstrates that E7e17_5 has larger $E_{B}$ and $E_{K}$ than in the weak-field runs.
+2 demonstrates that H7e17_5 has larger $E_{B}$ and $E_{K}$ than in the weak-field runs.
 $E_{K}$ follows the same qualitative trend as the other runs, with a rapid rise followed by a peak and slow decay in time (possibly with small oscillations).
-Unlike the other runs, however, $E_{B}$ immediately decays, while it remains nearly constant in E1e17_5 and grows for some time in both E1e16_5 and E6e15_5.
+Unlike the other runs, however, $E_{B}$ immediately decays, while it remains nearly constant in H1e17_5 and grows for some time in both H1e16_5 and H6e15_5.
 There is also a crossover between $E_{B}$ and $E_{K}$ at the same approximate time emergence occurs, though, as will be discussed later, this may simply be coincidence.
 III.2 Effects of density perturbation
-We now consider the effects of the perturbation by analyzing the non-perturbed cases, beginning with E1e17_0.
-Since E1e17_5, which has a density perturbation, does not emerge, it is not surprising that E1e17_0 does not, either, and instead rises only for a short time before stalling.
+We now consider the effects of the perturbation by analyzing the non-perturbed cases and the $k=9$ cases, beginning with H1e17_0.
+Since H1e17_5, which has a density perturbation, does not emerge, it is not surprising that H1e17_0 does not, either, and instead rises only for a short time before stalling.
 As shown in Fig.
-8, $E_{K}$ peaks at ${\sim}1/3$ its value in E1e17_5, and it rapidly decays by nearly an order of magnitude in ${\sim}0.3~{}\textrm{ms}$.
-Additionally, though the evolution of $E_{B}$ in E1e17_0 is initially similar to E1e17_5, it does not experience the same slow rise, but instead slowly decays over time.
-Figure 8: Like Fig. 2, but comparing the cases with no perturbation to their perturbed counterparts. The dotted vertical lines mark the emergence times for H7e17_0 and H7e17_5.
-Comparing the E7e17_0 case to E7e17_5, we find that the perturbation has very little effect on the dynamics of the system.
+8, $E_{K}$ peaks at ${\sim}1/3$ its value in H1e17_5, and it rapidly decays by nearly an order of magnitude in ${\sim}0.3~\textrm{ms}$.
+Additionally, though the evolution of $E_{B}$ in H1e17_0 is initially similar to H1e17_5, it does not experience the same slow rise, but instead slowly decays over time.
+Figure 8: Like Fig. 2, but comparing the cases with different perturbations ($A=0$, $k=9$, and $k=18$). The dotted vertical lines mark the emergence times for H7e17_0, H7e17_5, and H7e17_5_k9.
+Comparing the H7e17_0 case to H7e17_5, we find that the perturbation has very little effect on the dynamics of the system.
 This is consistent with our initial conditions, as the change in density from our equilibrium conditions on the flux tube ($\Delta\rho/\rho_{0}\approx 0.6$) is much larger than the change in density from the perturbation ($A=0.05$).
@@
 Fig.
-7 shows that emergence in E7e17_0 is slightly delayed to $t_{\mathrm{em}}\approx 0.224$ and slightly slower at $v_{\mathrm{em}}\approx 0.034c$ compared to E7e17_5, though the entrained mass is slightly larger.
+7 shows that emergence in H7e17_0 is slightly delayed to $t_{\mathrm{em}}\approx 0.224$ and slightly slower at $v_{\mathrm{em}}\approx 0.034c$ compared to H7e17_5, though the entrained mass is slightly larger.
 Nevertheless, the mass again rapidly decays with distance, suggesting that the amount of unbound ejecta is small.
+H1e17_5_k9 demonstrates sustained growth in the kinetic energy compared to H1e17_5 due and therefore rises further, which at first glance may suggest that $k=9$ may be closer to the critical wavenumber than $k=18$.
+However, the flux tube still fails to emerge, and the kinetic energy never exceeds what is observed in Fig.
+2 for H0_0_5.
+At early times, the growth rate is nearly identical for both perturbation modes, so it seems that the differences in evolution are not necessarily because $k=9$ initially excites the Parker instability but cannot overcome the stratification of the system near the surface; rather, it seems that the dynamics of both $k=9$ and $k=18$ are dominated by the hydrodynamic perturbation, and $k=9$ suppresses it less due to reduced magnetic tension.
+On the other hand, the H7e17_5_k9 case emerges at $t_{\mathrm{em}}\approx 0.194~\textrm{ms}$ with a velocity of $v_{\mathrm{em}}\approx 0.048c$, somewhat earlier than H7e17_5, and there is evidence that the kinetic energy is marginally larger.
+When comparing the kinetic energy evolution at early times for perturbed cases to H7e17_0, we find a maximum relative difference of ${\sim}9\%$ for H7e17_5_k9 compared to ${\sim}2\%$ for H7e17_5.
+This might suggest the $k=9$ perturbation grows faster, but we note that this difference does not demonstrate clear monotonic growth, so we claim with any confidence that this is more than a numerical effect.
+Additionally, assuming $E_{K}\propto t^{p}$ for some growth scale $p$, we do not see any meaningful difference in $p$ between the different perturbations.
+Fig.
+9 indicates that $d\log E_{K}/d\log t$ is effectively identical for all three cases, even at early times.
+Therefore, the growth timescale for the Parker instability must be much longer than the emergence timescale.
+This is not truly surprising, as the Parker instability does not usually dominate over other buoyant instabilities until it becomes nonlinear (see, e.g., Refs.
+Matsumoto et al.
+(1993); Nozawa (2005)).
+Figure 9: The growth rate $d\log E_{K}/d\log t$ as a function of time for H7e17_0, H7e17_5, and H7e17_5_k9.
 III.3 Effects of flux tube size
@@
 However, it is harder to distinguish the results from surface oscillations, as seen in the top half of Fig.
-9.
+10.
 This contamination also affects the additional integration surfaces (not shown), so the velocity at emergence is significantly more uncertain.
@@
 The energetics for H7e17_5_B are shown in the bottom plot of Fig.
-9.
+10.
 Qualitatively, the energetics are similar to H7e17_5, including the presence of a crossover around the time of emergence.
@@
 According to Fig.
-10, the energetics are qualitatively similar between the different resolutions, though the lower resolutions see $E_{B}$ decay faster and less growth in $E_{K}$, consistent with additional diffusion in the flux tube.
+11, the energetics are qualitatively similar between the different resolutions, though the lower resolutions see $E_{B}$ decay faster and less growth in $E_{K}$, consistent with additional diffusion in the flux tube.
 The crossover point between the two exists in all three runs, including the LR run that does not emerge, and seems to move earlier in time as the resolution decreases.
@@
 The mass flux increases monotonically with resolution except for the VLR case (see Fig.
-9), which appears to push out noticeably more mass than any other resolution.
+10), which appears to push out noticeably more mass than any other resolution.
 However, as in H7e17_5, the integrated mass decreases rapidly across the other integration surfaces (not shown), even for VLR, suggesting that it is still bound.
-There is little consistency in the merger time, and the effects of surface oscillations make it difficult to determine exactly where emergence occurs, particularly for the LR case.
+There is little consistency in the emergence time, and the effects of surface oscillations make it difficult to determine exactly where emergence occurs, particularly for the LR case.
 Though the flux tube is well resolved for the LR, MR, and HR cases (and arguably adequately resolved for the VLR case), it is likely that we are not in a completely convergent regime, even if we ignore the measured emergence times.
@@
 24, initially exceeded $50\%$ of the gravitational force on the flux tube in some of our tests.
-Alternate field configurations may be less susceptible to this tension force, but numerical simulations suggest that post-merger fields, independent of their initial (realistic) conditions, are dominated by a large-scale toroidal component [14, 40], which limits the possibilities for alternate field structures.
+Alternate field configurations may be less susceptible to this tension force, but numerical simulations suggest that post-merger fields, independent of their initial (realistic) conditions, are dominated by a large-scale toroidal component Aguilera-Miret et al.
+(2022); Gutiérrez et al.
+(2025), which limits the possibilities for alternate field structures.
 Even if this were not the case, this likely is only important for fields of ${\sim}10^{17}~\textrm{G}$; we observe that the evolution of the flux tube at weaker field strengths around $10^{16}~\textrm{G}$ are dominated by ordinary hydrodynamic buoyancy.
 Therefore, it seems unlikely that altering the field configuration is enough to help these features to emerge.
-We do not consider the effects of rotation in our model, but we remind the reader that rotation tends to inhibit convective processes, not enhance them [41, 42, 12].
-This result may initially seem at odds with works suggesting that strong magnetically driven outflows may develop from a BNS remnant, such as Most and Quataert [9] and Combi and Siegel [10].
-This is not necessarily the case; the demonstrated flare-like features and outflows in Most and Quataert [9] originate from strong magnetic fields exceeding $10^{17}~{}\textrm{G}$.
-Similarly, the post-merger jet in Combi and Siegel [10] develops in a remnant with an average toroidal field strength nearing and isolated pockets exceeding $10^{17}~{}\textrm{G}$.
+As stated in Sec.
+II.1, a rotating model is also unlikely to improve emergence, as rotation tends to inhibit convective processes, not enhance them Acheson and Gibbons (1978); Spruit (1999); Gao et al.
+(2026).
+This result may initially seem at odds with works suggesting that strong magnetically driven outflows may develop from a BNS remnant, such as Most and Quataert (2023) and Combi and Siegel (2023).
+This is not necessarily the case; the demonstrated flare-like features and outflows in Most and Quataert (2023) originate from strong magnetic fields exceeding $10^{17}~\textrm{G}$.
+Similarly, the post-merger jet in Combi and Siegel (2023) develops in a remnant with an average toroidal field strength nearing and isolated pockets exceeding $10^{17}~\textrm{G}$.
 However, we again stress that such field strengths are atypically large compared to observation of magnetars and what is expected in realistic merger remnants.
@@
 Musolino et al.
-[43], for example, suggest that, even with an extremely large magnetic field, these outflows may instead originate from the disk.
+(2025), for example, suggest that, even with an extremely large magnetic field, these outflows may instead originate from the disk.
 Similar to alternate configurations, buoyant instabilities like the Parker instability are likely only important for a small range of field strengths in excess of $10^{17}~\textrm{G}$.
 Despite a strong longitudinal perturbation, the Parker instability does not appear to develop in H1e17_5; rather, the flux tube oscillates.
-This agrees with results from Ref.
-[43], for instance, which suggest that the post-merger disk can be unstable to the Parker instability, but the remnant itself should be stable, even for a very large magnetic field.
+This agrees with results from Refs.
+Musolino et al.
+(2025); Jiang et al.
+(2025), for instance, which suggest that the post-merger disk can be unstable to the Parker instability, but the remnant itself should be stable, even for a very large magnetic field.
 Larger fields like H7e17_5 may be Parker-unstable, but the emergence time scale is so short that the instability cannot develop quickly enough to be dynamically important (as seen when compared with H7e17_0).
@@
 Though we see some differences in the emergence time (which may not be entirely accurate for H7e17_5_B), they are small enough that they would be insignificant in the context of a full merger simulation.
-This suggests that the timescales shown in global BNS runs with $\Delta x\lesssim 180~{}\mathrm{m}$ are likely reasonable despite the limited resolution.
+This suggests that the timescales shown in global BNS runs with $\Delta x\gtrsim 180~\mathrm{m}$ are likely reasonable despite the limited resolution.
 Secondary characteristics such as the amount of mass pushed out during emergence are more strongly affected by resolution.
@@
 In agreement with Musolino et al.
-[43], this suggests that the strong magnetically-driven outflows needed for BNS remnants to power relativistic jets may originate in the disk rather than the remnant.
+(2025), this suggests that the strong magnetically-driven outflows needed for BNS remnants to power relativistic jets may originate in the disk rather than the remnant.
 Though our work suggests emergence itself is somewhat independent of resolution so long as the large-scale magnetic field is well resolved, secondary features like the mass flux may be more sensitive to resolution.
@@
 We use 100 profiles of the $z$-axis equally spaced in time from the last $4.7$ ms of the remnant studied in Ref.
-[11] (${\sim}100~{}\mathrm{ms}$ in duration).
+Radice and Bernuzzi (2023) (${\sim}100~\mathrm{ms}$ in duration).
 For fits of $Y_{l}$, we compute $Y_{l}=Y_{e}+Y_{\nu_{e}}-Y_{\bar{\nu}_{e}}$, where $Y_{e}$, $Y_{\nu_{e}}$, and $Y_{\bar{\nu}_{e}}$ are the electron, electron neutrino, and electron antineutrino fractions, respectively.
 These fits are shown in the left half of Fig.
-11.
+12.
 This profile is extrapolated to high density based on these fits.
@@
 Figure 12: (Left) Fits of $T$ and $Y_{l}$ as functions of $\rho$ to the remnant from Ref.
-[11].
+Radice and Bernuzzi (2023).
 Black lines indicate profiles along the z-axis from the last $4.7$ msof the remnant, the red line marks the fit to this line.
@@
 Appendix B Relativistic Magnetic Buoyancy in the TOV Spacetime
-Note: for this section, we consider geometrized units with $G=c=k_{b}=1$ and adopt the Heaviside-Lorentz convention for the magnetic field; that is, $B^{i}_{\mathrm{HL}}=B^{i}_{\mathrm{CGS}}/\sqrt{4\pi}$.
+Note: for this section, we consider geometrized units with $G=c=1$ and adopt the Heaviside-Lorentz convention for the magnetic field; that is, $B^{i}_{\mathrm{HL}}=B^{i}_{\mathrm{CGS}}/\sqrt{4\pi}$.
 We also use the Einstein summation convention for paired indices, e.g., $a^{\mu}b_{\mu}\equiv\sum_{\mu}a^{\mu}b_{\mu}$.
-Consider the relativistic generalization of Eq. 1 [44]:
+Consider the relativistic generalization of Eq. 1 Lichnerowicz (1994):
 $\left(e+P+b^{2}\right)\frac{du^{\alpha}}{d\tau}+\mathcal{P}^{\alpha}_{\nu}g^{\mu\nu}\partial_{\mu}\left(P+\frac{b^{2}}{2}\right)-\mathcal{P}^{\alpha}_{\nu}\nabla_{\mu}\left(b^{\mu}b^{\nu}\right)=-\left(e+P+b^{2}\right)\Gamma^{\alpha}_{\mu\nu}u^{\mu}u^{\nu},$ (17)
@@
 24 is often a weaker constraint than its Newtonian counterpart.
-A massive remnant may have $m(r)\sim 2~{}\textrm{M}_{\odot}$ near a shear layer located at $r\sim 2~{}\textrm{M}_{\odot}$, suggesting that $(m(r)+4\pi r^{3}P_{0})/(r-2m(r))$ can be more than a factor of $2$ larger than $m(r)/r$.
+A massive remnant may have $m(r)\sim 2~\textrm{M}_{\odot}$ near a shear layer located at $r\sim 7~G\textrm{M}_{\odot}/c^{2}\approx 10.3~\mathrm{km}$, suggesting that $(m(r)+4\pi r^{3}P_{0})/(r-2m(r))$ can be more than a factor of $2$ larger than $m(r)/r$.
 The precise behavior of the term $\Delta e-\Delta P$ depends on the actual EOS, but a sufficient condition to ensure that $\Delta e-\Delta P>\Delta\rho$ is that $d(\rho\epsilon)/dP>1$.